### v1

# SoK: Payment Channel Networks

Abstract
Payment Channel Networks (PCNs) have been proposed as an alternative solution to the scalability, throughput, and cost overhead associated with on-chain transactions. By facilitating offchain execution of transactions, PCNs significantly reduce the burden on the blockchain, leading to faster transaction processing, reduced transaction fees, and enhanced privacy. Despite these advantages, the current research in PCNs presents a variety of research challenges that require further exploration.
In this paper, we survey the recent work in several aspects of PCNs, such as pathfinding and routing, virtual channels, state channels, payment channel hubs and rebalancing. This survey aims to provide the reader with a detailed understanding of the current state-of-the-art in PCN research, highlighting a few important advancements. Additionally, we highlight the various unresolved issues in the area of PCN research.
Specifically, this paper seeks to answer the following crucial question: What are the various interesting and non-trivial challenges in PCN research that require immediate attention from the academic and research community? By addressing this question, we aim to identify the most pressing problems and future research directions that interested readers can immediately work on. Through this analysis, we hope to inspire researchers and practitioners to tackle these challenges to make PCNs more secure and versatile.
Index Terms:
Payment channel network, Blockchain, Layer-2
I Introduction
(a) Payment channel opening
(b) Alice sending 30 coins
(c) Local balances get updated
(d) Posting final balances to blockchain
Figure 1: Payment channel operations
Cryptocurrencies and cryptocurrency based transactions have become increasing popular. Currently, the total market value of all cryptocurrencies in use has surpassed 2.5 Trillion USD. The cyrptocurrency market is increasing at a rate of $\approx$ 8.00% every year [1]. This rise in popularity can be attributed to the following reasons: 1) cryptocurrency transactions can be carried without the presence of a trusted entity. Fiat currency based transactions on the other hand, require the presence of a trusted financial organization such as a bank. 2) cryptocurrency transactions do not subject the user to any limits on the number and type of transactions. Fiat currency based transactions are limited in their amount and number, depending on several factors such as the currency, geographical location, etc. [2].
Each transaction posted to the Bitcoin blockchain takes around 7 seconds to be validated [3, 4]. The procedure of validation involves verifying that the transaction posted to the blockchain contains all the required fields and the signature of the user creating the transaction tuple is valid. Once the validation procedure is successfully completed, the transaction is included in a block that would be mined on the blockchain at some time in the future. The process of mining the block successfully takes $\approx$ 2 hours [5] (as of June $13^{th}$ 2024). This delay in the transactions and blocks getting confirmed is termed as the blockchain scalability problem [6]. In contrast, Visa, a company which globally processes transactions using fiat currencies, processes around 24,000 transactions per second [7]. Due to the delay in transaction processing caused by the blockchain scalability problem, blockchain-based transactions cannot process payments instantaneously.
As an alternative to processing transactions by posting to the blockchain, payment channels have been proposed. Two parties with the intent of processing payments between them open a payment channel by creating a transaction tuple called the funding transaction. This funding transaction contains the initial deposits from both the parties. These initial deposits are also called as the initial balances of the parties in the payment channel. The sum aggregate of these initial balances is called as the channel capacity. The funding transaction contains the signatures of both the parties involved in the payment channel making it a 2-2 multi signature transaction. This means that the funds in the funding transaction cannot be spent without the signatures of both the parties. This funding transaction is validated and included in a block. Once this block has been successfully mined and confirmed on the blockchain, the payment channel is opened between the two parties. The two parties can now be involved in an unlimited number of transactions with each other as long as the amount of a single transaction does not exceed their local balances.
An example of a payment channel is given in Figure 1(a). Two users (also called nodes) Alice and Celia deposit 50 coins each into a 2-2 multi signature transaction. This transaction is posted to the blockchain, upon which it is validated and included in a block. The block is mined and confirmed, at which time a payment channel is said to open between Alice and Celia. The sum aggregate of the individual balances of Alice and Celia, which is the channel capacity, is 100 coins. Alice making a payment of 20 coins to Celia is shown in the Figure 1(b) and the updated balances of Alice and Celia are shown in Figure 1(c). After the payment has been made, the channel capacity between Alice and Celia still remains constant at 100 coins. In this manner, Alice and Celia can be involved in an unlimited number of payments between each other. Each payment made in the payment channel in an off-chain manner contains the signatures of both Alice and Celia. When either Alice or Celia decide to close the payment channel, they post their final balances to the blockchain and the payment channel is closed as shown in Figure 1(d).
For each off-chain transaction in the payment channel, both the parties involved in the payment channel create a commitment. This commitment is essentially an agreement for the new balances signed by both the parties. Exchanging of commitments signifies that both the parties have agreed to the change in their respective balances. Each pair of commitments (for each transaction), contains a unique sequence number called the revocation sequence maturity number. For each new transaction made in the payment channel, the sequence number of the prior transaction is invalidated by revocation keys of both the parties. These revocation keys are created by both the parties before opening of the payment channel. If a malicious party in the payment channel broadcasts an older balance to the blockchain, the honest party in the channel has a certain time period during which it can contest this behavior on the blockchain. Before this time period expires, the honest party in the payment channel will broadcast the revocation of this old state signed by both the parties. The broadcasting of this revocation to the blockchain prevents the malicious party from stealing funds of the honest party.
The idea of a payment channel that exists between two parties can be extended to a number of nodes, creating a network of payment channels, called a payment channel networks or PCN. The advantage of PCNs is, they enable users that are not connected by a direct payment channel to make payments between each other in an off-chain manner. An example PCN is shown in Figure 2. In the figure, consider Alice who intends to make a payment to Hector, with whom she does not share a payment channel. The naive way to process this transaction would for Alice to open a payment channel with Hector, which involves Alice making an expensive blockchain write for the channel opening. Each payment channel opening costs 2.4 USD for blockchain writes [8, 9, 10]. If Alice intends to send an amount of 1 coin to Hector, it may not be economical for her to open a direct payment channel. Alice can make use of the PCN and make a payment to Hector by forwarding the payment along the path Alice $\rightarrow$ Celia $\rightarrow$ Michael $\rightarrow$ Rajiv $\rightarrow$ Charlie $\rightarrow$ Garcia $\rightarrow$ Hector. This process of using intermediate nodes in a PCN to forward to the payment to the intended destination is called as routing in payment channel networks.
Figure 2: Payment channel network
Motivation and timeliness of PCN research: A significant advantage of PCNs is their capability to facilitate micro-payment transactions, with minimum amounts as low as $10^{-7}$ BTC [11]. In contrast, the average transaction cost for a single on-chain transaction on the BTC blockchain is approximately $4.612$ USD as of June 2024 [12]  [^1] Opening a payment channel requires two inputs (for funds from both the parties) and one output (the funds are locked in a single 2-2 multi signature address). Where as making a payment by posting it to the blockchain requires one input (from the party that intends to make the payment) and two outputs (one to the receiver and one to the validator/miner for his fees/reward). This difference makes the transaction cost associated with opening a payment channel cheaper than making a payment using a blockchain write., regardless of the transaction amount. This cost can be avoided by using the off-chain PCNs, which incurs no additional fees. Additionally, transactions on the BTC blockchain take around 2 hours to be confirmed as of June 13, 2024 [5], whereas PCNs can process transactions instantly. An example of a real-world PCN is the Lightning Network ($\mathsf{LN}$) on the Bitcoin blockchain [13], which has a 24-hour trading volume of \$63,200 and a market capitalization of \$7 million [14] as of June 2024. These figures reflect $\mathsf{LN}$’s size and growth.
Contributions:
1) We qualitatively compare the recent work in several aspects of PCNs, viz. pathfinding and routing, virtual channels, state channels, payment channel hubs and rebalancing using several relevant properties (metrics) along with providing a reasoning why these metrics have been chosen for comparison.
2) In this paper, we point out the open problems in all the areas that we survey and we also discuss why solving those problems is a hard research challenge.
Outline: In Section II, we start with describing the concept of pathfinding and routing in PCNs and qualitatively compare work published in that area. In Section III, we describe virtual channels which have been proposed to address issues with multi-hop routing in PCNs, and compare work in this area. In Section III, we describe and compare state channel protocols, which are a more generalized adaptation of virtual channels and can facilitate the execution of arbitrary applications between nodes in the PCN (not just payments). In Section V, we cover payment channel hubs, which are similar to virtual channels and state channels, but facilitate a different use case for payments in PCNs. All of the aforementioned PCN mechanisms consume one common resource: the local balance of a node in a PCN. In Section IV, we discuss rebalancing, which addresses the important function of replenishing channel funds in the PCN. In Section VII, we present the reader with the current research gaps in all of the aforementioned areas and in Section IX, we conclude the paper.
Prior work: Prior works by Khojasteh et al. [15] and Erdin et al. [16] survey the work done only in the area of pathfinding and routing protocols and their privacy aspects in PCNs. Whereas, in this paper we cover the entire spectrum of PCN research: rebalancing, virtual channels and state channels, pathfinding and routing. Apart from this [15, 16] do not provide any information about the open problems in PCNs, which we do in our work. The SoK by Gudgeon et al. [17], surveys several layer-2 protocols, whereas, we focus exclusively on PCNs. Furthermore [17] was published in 2020 and does not cover most of the recent work published in PCNs.
TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs. A $\checkmark$ implies that the protocol satisfies the property and an implies that the routing protocol does not satsify the property
Protocols Concurrency Privacy Topology privacy Avoids source routing Decentralized Atomicity Disjoint graphs Fees
MPCN-RP [18] $\checkmark$ $\checkmark$ (flat)
Eckey et al [19] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Vein [20] $\checkmark$ (dynamic)
Auto tune [21] $\checkmark$ (flat)
Kadry et al. [22]
FSTR [23]
SilentWhispers [24] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Blanc [25] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SpeedyMurmurs [26] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Spider [27]
Flash [28] $\checkmark$
Coinexpress [29] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Webflow [30] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Robustpay [31] $\checkmark$ $\checkmark$ (flat)
Robustpay+ [32] $\checkmark$ $\checkmark$ (flat)
RACED [33] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Cheapay [34] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (flat)
Auroch [35] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Yang et al. [36] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SPRITE [37] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
II Pathfinding and Routing
Motivation: One of the areas in PCNs which has garnered significant attention from the academic community is pathfinding and routing. It is defined as the process of finding a path comprising of several nodes from a sender to a receiver in a PCN along which the payment intended to be sent by the sender can be forwarded to the receiver and forwarding the payment along that path.
Intuitively, it may seem that well-known pathfinding and routing protocols from the networks domain can be easily applied to PCNs. Unfortunately, there are several problems with this.
Even though PCNs are decentralized and distributed peer-peer networks, pathfinding and routing protocols of traditional networks cannot be applied to solve the pathfinding and routing problem in PCNs, because of fundamental differences between intent and the network features: 1) Transfer of data and transfer of money: The intent behind traditional network’s pathfinding and routing protocols is to transfer data between connected devices. In PCNs, the intent to transfer money between users in a decentralized and distributed manner. 2) Change in bandwidth and change in monetary state: When data is transferred between connected devices using network pathfinding and routing protocols, the bandwidth, which is represented by the link weights between the nodes does not change. However, transfer of money between users in a payment channel permanently alters the state of the sender and receiver involved by subtracting and adding money from and to their link weights respectively. 3) Latency and routing fees: The cost of sending data in network pathfinding and routing protocols is measured in terms of latency, which is defined as the time delay experienced in transferring data from one device to another. The delay in data transfer is experienced by the receiver and the cost of the delay is usually paid by the sender. The cost of sending payments in a PCN is measured in terms of money, called as routing fees. This encourages greedy behavior and makes the PCN prone to several attack vectors. Due to these reasons, pathfinding and routing in PCNs is a fundamentally different and a hard research challenge.
The properties in Table I represent the fundamental principles of fiat currency transactions and on-chain payment mechanisms, which we want reflected in off-chain payments. These properties are generally agreed upon in the literature by several works such as [33, 35, 37, 25, 26, 24] as common evaluation metrics for pathfinding and routing protocols in PCNs. Fulfilling these properties while providing efficient pathfinding and routing is a non-trivial challenge, and necessitates the design of new pathfinding and routing protocols. Several elegant pathfinding and routing protocols have been proposed in the literature. In Table I, we present a qualitative comparison of these routing protocols with respect to the properties they achieve.
Ideal properties:
1) Concurrency: Concurrency is achieved when a pathfinding and a routing protocol enables the nodes to forward more than one payment simultaneously.
2) Privacy: Privacy is maintained when a node’s identity is known only to its immediate neighbors and not to the entire network.
3) Topology Privacy: Topology privacy is preserved when no single node has knowledge of the entire network topology.
4) Avoids Source Routing: Source routing is avoided when the sender does not determine the path to the receiver.
5) Decentralization: Decentralization is achieved when there are no centralized or trusted entities responsible for constructing paths for senders.
6) Atomicity: Atomicity is ensured when the payment is routed all the way from the sender to the receiver or the payment is not routed at all.
7) Disjoint Graph Applicability: A pathfinding and routing protocol is considered applicable to disjoint graphs if it functions even when the network graph consists of islands.
8) Fees: Routing fees is the amount a node charges for forwarding the payment to the next node along a path from the sender to receiver. This fees can be charged in two ways. Flat/fixed fees means that the fees charged for routing payments remains the same irrespective of the transaction amount being routed. If the fees charged by a node varies according to the transaction amount, it is referred to as dynamic fees, typically a percentage of the amount being transacted.
III Virtual channels
(a) Virtual channel example [38, 39]
(b) Multi-hop virtual channel (the dots between Jill and Alice depict several intermediate nodes)
(c) Recursive virtual channel
Figure 3: Different types of virtual channels
TABLE II: Comparison of Virtual Channel Protocols
Protocol BC Validity Fee Privacy Off-chain dispute resolution Recursive Multihop
Perun [38] TC Limited
Jourenko et al. [40] UTXO Limited $\checkmark$
Aumayr et al. [41] UTXO Limited $\checkmark$ (fixed)
Elmo [42] UTXO Unlimited $\checkmark$
Donner [43] UTXO Unlimited $\checkmark$ (time based) $\checkmark$ $\checkmark$
Jia et al. [44] UTXO and TC Limited $\checkmark$
Eckey et al. [45] TC Limited $\checkmark$
Generalized state channels [46] TC Limited $\checkmark$
Transactions in PCNs are routed from the sender to the receiver using a path of intermediate nodes. This pathfinding and routing mechanism requires the nodes along the payment path to be available for a transaction to be processed. Nodes, however, sometimes can choose to go offline or there can be network/service disconnections. This might cause the transaction to fail. Virtual channels, which are built on top of existing payment channels, provide a solution to this problem. Virtual channels facilitate payments between a pair of nodes using an intermediate node. The intermediate node has individual payment channels open with the other two nodes. The intermediary and the pair of nodes lock coins with each other in their respective payment channels to establish a virtual channel. Once the virtual channel has been established, the two nodes can be involved in a unlimited number of payments. These payments can be processed even if the intermediate node goes offline.
The first virtual channel construction, Perun, was proposed by Dziembowski et al. in [38]. Assume there are two users, Alice and Bob, along with an intermediary as depicted in Figure 3(a). Alice establishes a payment channel, $\alpha_{A}$, with the intermediary by locking $Y_{A}$ coins as collateral; the intermediary contributes $Y_{I}$ coins into this channel. Similarly, Bob locks $Z_{B}$ coins into his channel, $\alpha_{B}$, with the intermediary, which also deposits $Z_{I}$ coins from its possession. This marks the completion of the opening phase, effectively locking the coins belonging to Alice and Bob. Following this, Alice and Bob create a virtual channel to facilitate transactions between them. The virtual channel will include coins deducted from both Alice’s and Bob’s payment channels. Specifically, $X_{A}$ coins will be subtracted from $Y_{A}$ on Alice’s side and deposited into the virtual channel, while $X_{B}$ coins will be subtracted from $Z_{B}$ and will be deposited from Bob’s side. Now Alice and Bob can be involved in an unlimited number of payments without the intermediary being online.
The idea of a virtual channel between a pair of nodes involving a single intermediary/intermediate node has been extended to establishing a virtual channel recursively over several hops involving several intermediaries, leading to the construction of a recursive virtual channel [40, 42, 43]. An example of a recursive channel is shown in the Figure 3(c). In the figure Trent and Alice are a pair of nodes who intend to establish a virtual channel between them. They do in three steps as follows; 1) Trent and Hu construct a virtual channel $V_{1}$ with Rajiv as an intermediary. 2) Hu and Alice construct a virtual channel $V_{2}$ with Jill as an intermediary. 3) Trent and Alice establish a virtual channel between them $V_{3}$.
A multihop virtual channel is constructed by establishing a single virtual channel between a pair of users over a path comprising of several intermediate nodes. An example of multihop virtual channel is shown in the Figure 3(b). Rajiv and Alice establish a single virtual channel $V$ between them.
Motivation: We have already described how PCNs offer a solution to the blockchain scalability problem by facilitating instantaneous payments in an off-chain manner. PCNs however, have few limitations: 1) The path generated in a PCN to process a payment between a sender and a receiver usually contains one or several intermediate nodes. All the nodes along the path are required to be available to route the payment until the receiver. This may not be possible at all times since, nodes can be disconnected from the network due to various reasons such as power outages, DoS attacks, etc. In a virtual channel, once the coins of the intermediary(ies) are locked, it can choose to remain offline without affecting the transaction processing. 2) Each node along the path from the sender to the receiver charges a fees called as routing fees/forwarding fees for forwarding the payment to the next node along the path. The sender has to pay all the nodes their respective forwarding fee along the path. This increases the overall cost of routing the payment. Virtual channels provide a solution to this problem, since they do not charge fees on a per transaction basis. Instead, the fees is charged based on a time epoch basis (the time for which the coins of the intermediary/intermediaries are locked in the virtual channel) or a fixed fee can be agreed in advance by all the parties involved in the virtual channel construction. 3) The time taken to route a payment from a sender to a receiver is directly proportional to the path length. A longer path would result in an increase in the overall pahtfinding and routing time. Virtual channels do not process payments through a path comprising of several intermediate nodes, hence the overall transactions’ pathfinding and routing time does not depend on the number of intermediate nodes between the sender and the receiver.
Virtual channels should not add an unnecessary burden to users, and should mirror the operations of payment channels as closely as possible, with the added benefit of no on-chain transactions at all, while maintaining comparable security/privacy properties. We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
Ideal properties:
1) BC: BC is the blockchain on which the corresponding virtual channel protocol can be deployed. TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO based scripting mechanism such as Bitcoin. Ideally, a virtual channel protocol should be deployable on any blockchain without the need for any additional programming/scripting requirements. The virtual channel protocol should not require any modifications to the existing consensus mechanisms of the underlying blockchains.
2) Validity: This determines the validity of the virutal channel. A limited validity means that the virtual channel is valid for a predetermined time period (which is decided by parties involved in the virtual channel). Unlimited validity means that the virtual channel can stay open until the parties involved decide to initiate closing. This provides better and efficient transaction processing between users who intend to have frequent payments made between each other.
3) Fee: This metric determines if the virtual channel takes into account the fees charged by the intermediate node(s) involved in the channel’s establishment. The fixed fee model implies that a predetermined, fixed fee is paid to the intermediate node(s) which is agreed upon by all the parties in the virtual channel and the time based fee model implies that a fixed fee is paid to the intermediary which depends on the time for which the virtual channel stays open. Having an well-defined fee structure will motivate the intermediary/intermediaries to participate in virtual channel creation.
4) Privacy: Similar to payment channels, we want privacy of nodes for virtual channels, which is said to be achieved when the identities of the nodes in the virtual channel protocols are not known to anyone except for their immediate neighbors.
5) Offchain dispute resolution: This metric determines if the transaction disputes in a virtual channel require a blockchain write.
6) Recursive: A virtual channel is said to be recursive or hierarchical if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing virtual channels on top of existing virtual channels from sender to receiver. We give an example of a recursive virtual channel in Figure 3(c).
7) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing a single virtual channel from the sender to the receiver, without establishing virtual channel between any pair of intermediate nodes along the path. We give an example of a recursive virtual channel in Figure 3(b).
(a) State channel opening
(b) Chess game using state channel
(c) State channel close
Figure 4: State channel example with a chess game decentralized application (DApp)
TABLE III: Comparison of State Channel Protocols
Protocol BC General/purpose state channels Security Graceful exit Off-chain dispute resolution
Pisa [47] TC $\checkmark$
Hydra [48] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Sprites [49] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Aumayr et al. [50] UTXO, TC $\checkmark$ $\checkmark$
Origami [51] TC $\checkmark$ $\checkmark$
Forcemove [52] TC $\checkmark$
IV State channels
A state channel is an off-chain protocol that can facilitate execution of an arbitrary decentralized application, also called as DApp (such as a two player game) between two users in a decentralized and distributed network. An example of a state channel is given in Figure 4(a). Let us consider two users Alice and Celia who intend to play a game of chess. They initially interact with the chess game application deployed on the blockchain by a third party service provider. Both Alice and Celia deposit 30 coins each from their possession into a 2-2 multi signature transaction (step 1). Upon locking the funds, the chess game is instantiated between Alice and Celia by the DApp (step 2). Alice and Celia engage in a series of moves for the chess game (step 3, step 4) and let us assume Alice eventually wins (step 5) as shown in Figure 4(b). Alice gets paid 15 coins as the prize money in Figure 4(c). In this game, Alice and Celia countersign each other’s moves until there is a winner or a draw. In the event of malicious activity by one party (such as undoing a prior move), the honest party will broadcast all the moves to the blockchain for dispute resolution.
Motivation: State channels allow the already existing payment channels to facilitate execution of arbitrary applications such as games, e-commerce, etc. This leads to the PCN becoming more versatile. State channels should not add an unnecessary burden to users in terms of programming/scripting requirements, with the added benefit of facilitating the execution of any decentralized application between them, while maintaining comparable security properties and should ideally be able to resolve disputes without accessing the blockchain. We now give the properties desired from an ideal state channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of various state channel protocols in Table III. We describe the metrics used for comparison below.
Ideal properties:
1) Blockchain: This is the blockchain on which the corresponding state channel protocol can be deployed. TC is a blockchain that supports a Turing complete programming language and UTXO if a blockchain that supports a UTXO based scripting.
2) General purpose state channels: This is the ability of the state channel to facilitate the execution of any application in an off-chain manner by using a single application-agnostic instance of state channel on the blockchain.
3) Security: Security is achieved by a state channel protocol if no honest party loses funds because of malicious behavior by other parties in the protocol.
4) Graceful exit: A state channel protocol is said to employ graceful exit if it has clear and well-defined mechanisms for nodes joining or leaving the state channel.
5) Off-chain dispute resolution: This metric indicates if a dispute that occurs between the parties involved in the state channel can be resolved without accessing the blockchain.
V Tumblers
TABLE IV: Comparison of Tumblers
Protocol BC
Relationship
anonymity
Privacy against
aborts
Time epochs
Dynamic
corruption
Atomicity
Value
privacy
Variable
amount
BOLT [53] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Nocust [54] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MIXCT [55] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
A2L [56]
UTXO and
TC
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Blindhub [57]
UTXO and
TC
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Teechain [58] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Accio [59] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Turbo [60] TC $\checkmark$ $\checkmark$
Boros [61] TC $\checkmark$ $\checkmark$ $\checkmark$
Tumblebit [62] UTXO $\checkmark$ $\checkmark$ $\checkmark$
Motivation: A payment channel hub is a layer-2 mechanism that facilitates transaction processing between a sender/receiver pair that do not have a payment channel open between them. The intermediary which facilitates payments is called a tumbler. The tumbler has multiple payment channels open with several nodes in the network to facilitate payments. The hubs can potentially be useful in facilitating micropayments between a buyer and seller, without either of them needing to open a payment channel with each other.
Payment channel hubs can be classified into two types: on-chain and off-chain. Early hubs, also called hubs or mixers, were on-chain [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80], but they all suffered from scalability issues due to having to post each transaction on the blockchain. This led to the development of offchain payment channel hubs for specific blockchains, e.g., Bolt [53] for Zcash, Nocust [54] and MixCT [55] for Ethereum. The most general-purpose payment channel hubs are Tumblebit [62], $A^{2}L$ [56], and Blindhub [57].
A payment channel hub should should be able to facilitate anonymous payments between a pair of sender/receiver with the added benefit of ensuring that the payment is sent to the receiver or it does not go through at all, while maintaining privacy properties. We now give the properties desired from an ideal payment channel hub, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of off-chain payment channel hubs in the Table IV. We describe the comparison metrics below:
Ideal properties:
1) Blockchain: BC is the blockchain on which the corresponding Tumbler protocol can be deployed. TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO based scripting mechanism such as Bitcoin.
2) Relationship anonymity: It ensures that the relationship between a sender and a receiver for a given transaction should not be known to the tumbler.
3) Privacy against aborts: Tumbler should not be able to deduce the identities of a sender/receiver in case of a transaction abort, regardless of which party is responsible for the abort.
4) Time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch. A payment channel hub that supports transaction processing based on time epochs is not efficient. Since nodes would ideally need to wait until a new time epoch begins to process payments again. The duration between time epochs depends on several factors such as the unit of the time epoch (absolute time or relative time measured in terms of block height), number of users that will most likely involved in the next time epoch, etc.
5) Dynamic corruption: The payment channel hub should be able to handle the deviation of any party from the protocol at any point during its execution. i,e the set of adversaries should not be static during the protocol execution.
6) Atomicity: Atomicity ensured when a payment is routed all the way from the sender to the receiver or the payment is routed at all.
7) Value privacy: Value privacy is ensured by a payment channel hub, when, given a transaction amount, the tumbler cannot link it to a sender/receiver pair.
8) Variable amount: The tumbler should be able to process transactions of any value.
TABLE V: Comparison of Rebalancing Protocols
Protocol BC Trusted entity Graph compatability Privacy
Revive [81] TC $\checkmark$ (C) Cycles only
Hide & Seek [82] UTXO $\checkmark$ (D) Cycles only $\checkmark$
Subramanian et al. [83] TC and UTXO Agnostic $\checkmark$
Cycle [84] TC Cycles only $\checkmark$
Rebal [85] UTXO Cycles only $\checkmark$
Shaduf [86] TC Agnostic $\checkmark$
Chen [87] et al. UTXO Cycles
Musketeer [88] UTXO Cycles
VI Rebalancing
Motivation:Rebalancing in payment channel networks involves redistributing the balance within an existing payment channel or across multiple channels to maintain or improve the network’s liquidity and usability. This is done to ensure that both parties in the channel have sufficient funds on their respective sides to send and receive payments. Rebalancing is crucial to payment channel networks. If the link weights of nodes become zero/depleted as a result of being involved in transactions, it will prevent them from taking part in further payments until those link weights are replenished. Such nodes are called as dormant nodes in the network. A PCN which has a large number of such dormant nodes experiences an overall reduction in transaction throughput and an overall reduction in liquidity. Hence, the link weights of the nodes in the PCN need to be rebalanced to prevent failure of transactions due to lack of liquidity. A rebalancing protocol should be be able to replenish the link weight of dormant nodes in the PCN irrespective of the PCNs topology, while not having to employ any trusted entity and should not compromise privacy of nodes. We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
We give a qualitative comparison of rebalancing protocols in Table V. We describe the metrics as follows:
Ideal properties:
1) Blockchain: This is the blockchain on which the corresponding rebalancing protocol can be deployed. As it can be inferred from the Table, the protocols can either be deployed on the Ethereum or the Bitcoin blockchains.
2) Trusted entity: This metric determines the nature of trust required among parties to deploy the rebalancing protocol. A centralized trust (C) indicates the presence of a single trusted entity and distributed trust (D) indicates the distribution of trust across various entities.
3) Graph compatibility: This indicates the nature of the topology of the network on which a rebalancing protocol can be deployed. Cycles indicate that the rebalancing protocol can only deployed on cyclic graph and agnostic indicates that the rebalancing protocol can be deployed irrespective of the network topology.
4) Privacy: Privacy is said to be achieved when sensitive information such as the local balance and the identities of nodes are not known to anyone except for their immediate neighbors.
VII Research Gaps & Open Problems
In this section, we highlight the gaps in research published until now in the areas of pathfinding and routing, virtual channel construction, payment channel hubs, and state channels. This section provides an overview of the current research problems in these areas along with a discussion on why is developing a solution to these problems an interesting research challenge.
VII-A Gaps in pathfinding and routing for PCNs
We now identify the gaps in research in pathfinding and routing protocols for payment channel networks.
1) Super node liquidity validation in PCNs: In payment channel networks, pathfinding and routing, which is finding a path between the sender and the receiver that can route the amount specified by the sender has attracted significant attention from the research community. Several elegant and protocols have been proposed to solve this problem [28, 89, 21, 24, 25] . In a PCN, nodes in the network store information about other nodes and their payment channels in their local storage called gossip table/gossip network. The amount of information a node has in its gossip table is a function of the number of payment channels it has open in the network. It is also a function of how much information nodes in the PCN are willing to share with their neighbors (nodes with which they have open payment channels) as a part of the gossiping protocol. The amount of information a node can attain using gossiping may be limited due to the following reasons: a node choose to share information about the network partially, a node can also choose to be anonymous, or a node’s own information can sometimes be limited due to the lack of having open payment channels.
One of the main problems with the super nodes is that the sender has no way to know whether the super node possess the local balance (liquidity) to route the payment of the sender. The local balance of a super node or any node in a PCN is a private value and should not be known to any node except for its immediate neighbor. Currently, if a super node does not have enough liquidity to route the payment of a sender, the transaction fails and it has to be retried by the sender. In $\mathsf{LN}$, one of the most widely used PCN, this is a significant problem. Sometimes the sender might have to keep retrying for $\approx$ 1 hour to have a successful payment [90]. The main goal of PCNs is to facilitate instantaneous payments and these transaction retries render such payments almost impossible. It might greatly benefit the sender if the it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns in the network.
2) Channel verification in a PCN: To be a part of any PCN, nodes will open payment channels with other nodes in the network. Two nodes open a payment channel between them by posting a transaction to the blockchain. This transaction can be posted on the blockchain (as in UTXO based blockchains like BTC) or this transaction can be posted as a function call to an existing smart contract (as in blockchains that support smart contracts like ETH). In the most popular PCN, $\mathsf{LN}$, the channel verification procedure is very inefficient. It involves the node who intends to verify a channel to request the block in which the channel opening transaction has been included, verifying whether the transaction has been successfully executed by the validator/miner and finally verifying if the channel opening transaction corresponds to a 2-2 multi signature address on the blockchain. The verifier performing these steps is inefficient since all these steps will have already been performed by the miner. According to the current verification process any node would need to perform any these steps to all the payment channels its neighbors have open. Consider the following scenario: a node $j$ in the PCN has payment channels open with $m$ nodes in the PCN. Furthermore, assume that each neighbor opens $k$ new channels for each time epoch. For this example, assume that the time epoch has been set to 1 block. This would mean that the total number of channels the node $j$ has to verify a total of $m*k$ channels. This requires the node $j$ to access the blockchain a total of $m*k$ times to perform channel verification. Finding a way to do this without blockchain access and in a blockchain agnostic manner in a hard research challenge.
VII-B Gaps in Tumblers
We will now describe the research gaps in tumblers.
1) Privacy against aborts: In a given a transaction, the tumbler should not be able to correctly identify the sender/receiver pair involved. Currently, there is only one construction [59] that guarantees this. However, this construction works only for blockchains that support Turing complete programming languages. Currently, there is no such construction for UTXO based blockchains that guarantees privacy against aborts. This is hard because if the sender aborts a transaction, the tumbler will know which receiver did not receive the payment. This is hard because, a tumbler mixes multiple transactions together to ensure unlikability among payments. Also, a tumbler should enforce atomicity so that the payment either goes through completely or does not go through at all. In the case of an abort, an adversary who controls the tumbler might be able to deduce which receiver amongst the ones involved did not receive a payment and will be able to link the corresponding transaction amount to that receiver.
2) Time epochs: A tumbler should be able to facilitate payments without having a predefined time interval or epoch. Currently, the majority of tumbler constructions process payments in discrete epochs.To increase their anonymity sets. An anonymity set refers to a set of users (senders/receivers) whose identity is indistinguishable from each other. The larger is the size of this set, more transactions will be mixed together, increasing privacy of sender/receiver in a transaction. When the interval between two consecutive time epochs is sufficiently large, it will promote more users to take part in the transactions within a time epoch, which improves privacy. Not having time epochs, would mean the tumbler can process transactions at any instant, resulting in smaller anonymity sets and hence could possibly weaken privacy. The trade off is between privacy and flexibility. Processing transactions in discrete time epochs results in increased privacy and reduced flexibility. However, the tumbler’s ability to process transactions at any time instant might result in smaller anonymity sets, reducing privacy.
3) Dynamic corruption: A tumbler should ideally be resistant to dynamic corruption. This means the protocol should be resilient to parties deviating from their ideal execution during the beginning, progress, and ending of the protocol’s execution. Current tumbler constructions provide support for static corruption, which means that an adversary decides beforehand the parties they intend to corrupt and that cannot change during the protocol execution, which may be unrealistic from a practical standpoint. Security against dynamic corruption is hard to achieve because the tumbler mixes several transactions together to preserve anonymity. Since the tumbler should not be able to link transaction amount to a sender/receiver pair (ideally), if a party gets corrupted during the protocol’s execution, the tumbler cannot correctly map the corrupted party to a particular transaction. If the sender/receiver becomes malicious during the protocol’s execution, ideally, that transaction must be aborted to prevent financial loss to the honest party. However, the tumbler should not be able to link a transaction to a sender/receiver pair, making security against dynamic corruption a hard research problem.
VII-C Gaps in Virtual Channels
We will now describe the open problems in the area of virtual channels.
1) Blockchain agnostic: A virtual channel should ideally be blockchain agnostic, meaning it should be deployable regardless of the underlying scripting or programming language structure supported by the blockchain.
2) Well defined fee structure: The intermediary/intermediaries involved in the virtual channel construction additionally lock coins apart from the ones locked in the underlying payment channel. It is necessary to have a well-defined fee structure for this additional service. Currently, nodes get paid a routing fees/forwarding fees for every transaction they process. This fee is dependent on the transaction amount being processed. In the case of virtual channels, having a well-defined fee structure is difficult due to the following reasons: 1) The fee structure should take into account the amount of funds and the time for which these funds of the intermediary/intermediaries are locked in virtual channel. 2) It also needs to take into account the routing/forwarding fees an intermediary could have earned by not locking up the coins in the virtual channel. The fees for an intermediary should be ideally be more than the routing fees it would have earned in order to motivate nodes to be involved in virtual channel construction. Currently, no virtual channel construction has a well-defined fee structure.
3) Off-chain dispute resolution: Ideally a virtual channel protocol should provide the flexibility to resolve any disputes without accessing the blockchain. We have seen in Section I, that blockchain writes are expensive. Having an off-chain dispute resolution is a hard research challenge due to the following reasons; 1) Any form of layer-2 disputes are usually resolved on the blockchain. Since underlying consensus in blockchain usually guarantees resolution of a dispute raised within a specific time period. In an off-chain scenario, there is no such mechanism which guarantees this. None of the virtual channel constructions have a provision of off-chain dispute resolution.
4) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel(s) between any pair of intermediate nodes along the path. Currently, no VC construction has this capability. This is hard because, multihop virtual channel involves designing a mechanism for sequentially locking coins across a path comprising of several intermediate nodes. Furthermore, the construction of a multihop virtual channel should also ensure that any transaction does not effect the overall balance neutrality (no node apart from the receiver along the path gains or loses any money). Also, the goals of privacy, atomicity, offchain dispute resolution need to be enforced.
VII-D Gaps in state channel protocols
We will now describe the gaps in the area of state channel protocols.
1) General purpose state channels: An ideal blockchain should be able to facilitate deployment and execution of several decentralized applications without requiring the users to access the blockchain for any purpose other than opening and closing the state channel. Currently, there are two such constructions [49, 48] which incorporate code reusability into their design. These two constructions support UTXO based blockchains and have no provision for offchain dispute resolution.
2) Offchain dispute resolution: Current state channel protocols use an onchain transaction or a function call to the onchain smart contract in the case of dispute resolution. An ideal state channel protocol should be able to facilitate dispute resolutions in an offchain manner. This is a hard research challenge since disputes on the blockchain are usually resolved by the underlying consensus mechanisms. In an offchain scenario, there is no such consensus that guarantees transaction validity. Also, in an offchain protocol such as the state channel, nodes which are not part of the state channel protocol do not know any details of the protocol’s execution due to privacy concerns. Onchain transactions on the hand are publicly accessible.
VIII Network Statistics
(a) Minimum accepted value of HTLC payment (msat)
(b) Distribution of base fee (msat)
(c) Distribution of rate fee (msat)
(d) Distribution of node degree
(e) Distribution of HTLC expiration time
Figure 5: $\mathsf{LN}$ statistics
This section introduces a general analysis of the parameters used in $\mathsf{LN}$, extracted from $\mathsf{LN}$ gossip messages [91]. It offers views into user behavior and the implications of such behavior for network performance. Our analysis shows that the nodes prefer to choose the default value of the $\mathsf{LN}$ parameters. Using standard default values for these parameters allows nodes to operate smoothly together, reducing the chance of compatibility issues and enhancing the overall reliability and efficiency of the $\mathsf{LN}$. We evaluated the analysis on a snapshot of $\mathsf{LN}$ topology from November 27, 2021 [92, 93]. The snapshot includes 18,331 nodes and 80,918 channels with a mean channel capacity of 4056152.35 satoshi and a median channel capacity is 1000000 satoshi (this provides us information on how funds are distributed among $\mathsf{LN}$ channels).
Although not explicitly shown in this chart, nodes with multiple connections suggest the presence of fewer central nodes compared to the end-users nodes. Figure 5(a) shows that the majority of the nodes (87.9%) set the ’htlc_minimum_msat’ to a default value of 1000 msat, which indicates that the $\mathsf{LN}$ supports micropayment transactions. Figure 5(b) shows that ’fee_base_msat’ parameter across all of the nodes has the highest two values 1000 msat as a default value, and 0 msat as the minimum value, which indicates that most of the nodes do not set the fee_base_msat beyond the default configuration, and likely to support the network usage. Figure 5(c) shows the majority of the nodes (51.8%) in $\mathsf{LN}$ use various ’fee_rate_milli_msat’ values. A substantial portion of nodes (35.2%) set their fee rate to 1 msat to attract more transactions. Also, different nodes have different fee rate strategies, making $\mathsf{LN}$ a competitive environment. Figure 5(d) shows that the majority of the nodes (49.5%) have only one connection that indicates that in $\mathsf{LN}$, many end-users connected to fewer well-connected nodes. Moreover, 30% of the nodes have no payment channels, these could be new nodes added to the $\mathsf{LN}$ or they could be inactive nodes. Figure 5(e) shows that (81.5%) of nodes prefer to set the ’cltv_expiry_delta’ to the default value of 40 blocks, which is the number of the blocks a node can wait before losing its satoshis from $\mathsf{LN}$. Moreover, it helps to increase the flexibility of the $\mathsf{LN}$ to include more nodes along the payment path before reaching the ”locktime max’ parameter value (2016 blocks). According to the analysis, the nodes generally prefer to use the default parameter values of the lightning network. Using these default values can help in developing more effective routing protocols and optimizing routing fee mechanisms for the nodes.
IX Conclusion
In this paper, we qualitatively compared the recent work in various areas of PCN research: pathfinding and routing, virtual channels, payment channel hubs, state channels and rebalancing. We have also provided the gaps in research in these areas along with reasons why fulfilling those gaps is a hard research challenge. We hope that this paper motivates researchers to build robust protocols that address these gaps that would go a long way towards building out and developing the decentralized finance ecosystem.

### v2

@@
 Furthermore [17] was published in 2020 and does not cover most of the recent work published in PCNs.
-TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs. A $\checkmark$ implies that the protocol satisfies the property and an implies that the routing protocol does not satsify the property
+TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs.
 Protocols Concurrency Privacy Topology privacy Avoids source routing Decentralized Atomicity Disjoint graphs Fees
@@
 Cheapay [34] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (flat)
-Auroch [35] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Auroch [35] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (dynamic)
 Yang et al. [36] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
 Donner [43] UTXO Unlimited $\checkmark$ (time based) $\checkmark$ $\checkmark$
-Jia et al. [44] UTXO and TC Limited $\checkmark$
+Jia et al. [44] UTXO and TC Limited $\checkmark$ (fixed)
 Eckey et al. [45] TC Limited $\checkmark$
@@
 3) The time taken to route a payment from a sender to a receiver is directly proportional to the path length.
-A longer path would result in an increase in the overall pahtfinding and routing time.
+A longer path would result in an increase in the overall pathfinding and routing time.
 Virtual channels do not process payments through a path comprising of several intermediate nodes, hence the overall transactions’ pathfinding and routing time does not depend on the number of intermediate nodes between the sender and the receiver.
 Virtual channels should not add an unnecessary burden to users, and should mirror the operations of payment channels as closely as possible, with the added benefit of no on-chain transactions at all, while maintaining comparable security/privacy properties.
-We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
+We now give the properties desired from an ideal virtual channel, and compare the works in this area in Table II on the extent to which they achieve these properties.
 Ideal properties:
@@
 We give an example of a recursive virtual channel in Figure 3(c).
-7) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing a single virtual channel from the sender to the receiver, without establishing virtual channel between any pair of intermediate nodes along the path.
-We give an example of a recursive virtual channel in Figure 3(b).
+7) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel between any pair of intermediate nodes along the path.
+We give an example of a multihop virtual channel in Figure 3(b).
 (a) State channel opening
@@
 This leads to the PCN becoming more versatile.
-State channels should not add an unnecessary burden to users in terms of programming/scripting requirements, with the added benefit of facilitating the execution of any decentralized application between them, while maintaining comparable security properties and should ideally be able to resolve disputes without accessing the blockchain.
+State channels should be deployable on any blockchain regardless of the underlying programming/scripting requirements and should be able to facilitate the execution of any decentralized application between the parties involved, while maintaining comparable security properties (in comparison to payment channels) and should ideally be able to resolve disputes without accessing the blockchain.
 We now give the properties desired from an ideal state channel, and compare the works in this area on the extent to which they achieve these properties.
@@
 The most general-purpose payment channel hubs are Tumblebit [62], $A^{2}L$ [56], and Blindhub [57].
-A payment channel hub should should be able to facilitate anonymous payments between a pair of sender/receiver with the added benefit of ensuring that the payment is sent to the receiver or it does not go through at all, while maintaining privacy properties.
+A payment channel hub should be able to facilitate payments between a sender and receiver, who do not have a payment channel open between them such that the hub cannot link a given transaction amount to a particular sender/receiver pair.
+Furthermore, the payment channel hub should also guarantee the fundamental property of atomicity (ensuring that the payment is sent to the receiver or it does not go through at all).
 We now give the properties desired from an ideal payment channel hub, and compare the works in this area on the extent to which they achieve these properties.
@@
 4) Time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch.
-A payment channel hub that supports transaction processing based on time epochs is not efficient.
-Since nodes would ideally need to wait until a new time epoch begins to process payments again.
+A payment channel hub that supports transaction processing based on time epochs is not efficient because nodes cannot process payments between time epochs, which can be of arbitrary length.
 The duration between time epochs depends on several factors such as the unit of the time epoch (absolute time or relative time measured in terms of block height), number of users that will most likely involved in the next time epoch, etc.
@@
 We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
-We give a qualitative comparison of rebalancing protocols in Table V. We describe the metrics as follows:
+We give a qualitative comparison of rebalancing protocols in Table V.
+We describe the metrics as follows:
 Ideal properties:
@@
 We now identify the gaps in research in pathfinding and routing protocols for payment channel networks.
-1) Super node liquidity validation in PCNs: In payment channel networks, pathfinding and routing, which is finding a path between the sender and the receiver that can route the amount specified by the sender has attracted significant attention from the research community.
-Several elegant and protocols have been proposed to solve this problem [28, 89, 21, 24, 25] .
-In a PCN, nodes in the network store information about other nodes and their payment channels in their local storage called gossip table/gossip network.
-The amount of information a node has in its gossip table is a function of the number of payment channels it has open in the network.
-It is also a function of how much information nodes in the PCN are willing to share with their neighbors (nodes with which they have open payment channels) as a part of the gossiping protocol.
-The amount of information a node can attain using gossiping may be limited due to the following reasons: a node choose to share information about the network partially, a node can also choose to be anonymous, or a node’s own information can sometimes be limited due to the lack of having open payment channels.
+1) Super node liquidity validation in PCNs: A super node is a highly connected node (it has payment channels open with numerous other nodes in the network) and also has a high liquidity in those channels.
+These super nodes aid other nodes in pathfinding and routing of payments.
+These super nodes are called as a trampoline nodes in $\mathsf{LN}$ [89, 90, 91, 92].
+Several other works in the literature refer to super nodes as routing node, routing helper, landmark node, routers, etc [25, 24, 27, 37, 33].
 One of the main problems with the super nodes is that the sender has no way to know whether the super node possess the local balance (liquidity) to route the payment of the sender.
@@
 It might greatly benefit the sender if the it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns in the network.
+Achieving this is hard because the local balance a node has with its immediate neighbors in a PCN is considered private information and the sender has to know whether the super node possess enough liquidity without comprising privacy.
 2) Channel verification in a PCN: To be a part of any PCN, nodes will open payment channels with other nodes in the network.
@@
 The verifier performing these steps is inefficient since all these steps will have already been performed by the miner.
-According to the current verification process any node would need to perform any these steps to all the payment channels its neighbors have open.
+According to the current verification process, any node would need to perform all of these steps for each payment channel it has open in the network.
 Consider the following scenario: a node $j$ in the PCN has payment channels open with $m$ nodes in the PCN.
 Furthermore, assume that each neighbor opens $k$ new channels for each time epoch.
-For this example, assume that the time epoch has been set to 1 block.
-This would mean that the total number of channels the node $j$ has to verify a total of $m*k$ channels.
+For this example, assume the value of the time epoch is 1 block.
+This would mean that the total number of channels the node $j$ has to verify a total of $m*k$ channels for every block.
 This requires the node $j$ to access the blockchain a total of $m*k$ times to perform channel verification.
@@
 2) Time epochs: A tumbler should be able to facilitate payments without having a predefined time interval or epoch.
-Currently, the majority of tumbler constructions process payments in discrete epochs.To increase their anonymity sets.
+Currently, the majority of payment channel hub constructions process payments in discrete epochs to increase the size of their anonymity sets.
 An anonymity set refers to a set of users (senders/receivers) whose identity is indistinguishable from each other.
-The larger is the size of this set, more transactions will be mixed together, increasing privacy of sender/receiver in a transaction.
+The size of this set is proportional to the privacy of sender/receiver in the transaction.
 When the interval between two consecutive time epochs is sufficiently large, it will promote more users to take part in the transactions within a time epoch, which improves privacy.
-Not having time epochs, would mean the tumbler can process transactions at any instant, resulting in smaller anonymity sets and hence could possibly weaken privacy.
+Not having time epochs results in smaller anonymity sets and hence could possibly weaken privacy.
 The trade off is between privacy and flexibility.
-Processing transactions in discrete time epochs results in increased privacy and reduced flexibility.
-However, the tumbler’s ability to process transactions at any time instant might result in smaller anonymity sets, reducing privacy.
 3) Dynamic corruption: A tumbler should ideally be resistant to dynamic corruption.
 This means the protocol should be resilient to parties deviating from their ideal execution during the beginning, progress, and ending of the protocol’s execution.
-Current tumbler constructions provide support for static corruption, which means that an adversary decides beforehand the parties they intend to corrupt and that cannot change during the protocol execution, which may be unrealistic from a practical standpoint.
+Current tumbler constructions only provide support for static corruption, which may be unrealistic from a practical standpoint.
 Security against dynamic corruption is hard to achieve because the tumbler mixes several transactions together to preserve anonymity.
-Since the tumbler should not be able to link transaction amount to a sender/receiver pair (ideally), if a party gets corrupted during the protocol’s execution, the tumbler cannot correctly map the corrupted party to a particular transaction.
-If the sender/receiver becomes malicious during the protocol’s execution, ideally, that transaction must be aborted to prevent financial loss to the honest party.
+Since the tumbler should not be able to link transaction amount to a sender/receiver pair (ideally), if the sender/receiver becomes malicious during the protocol’s execution, ideally, that transaction must be aborted to prevent financial loss to the honest party.
 However, the tumbler should not be able to link a transaction to a sender/receiver pair, making security against dynamic corruption a hard research problem.

### v3

@@
 Abstract
-Payment Channel Networks (PCNs) have been proposed as an alternative solution to the scalability, throughput, and cost overhead associated with on-chain transactions.
+Payment Channel Networks (PCNs) have been proposed as an alternative solution to the scalability, throughput, and cost overhead problems associated with blockchain transactions.
 By facilitating offchain execution of transactions, PCNs significantly reduce the burden on the blockchain, leading to faster transaction processing, reduced transaction fees, and enhanced privacy.
-Despite these advantages, the current research in PCNs presents a variety of research challenges that require further exploration.
-In this paper, we survey the recent work in several aspects of PCNs, such as pathfinding and routing, virtual channels, state channels, payment channel hubs and rebalancing.
-This survey aims to provide the reader with a detailed understanding of the current state-of-the-art in PCN research, highlighting a few important advancements.
-Additionally, we highlight the various unresolved issues in the area of PCN research.
-Specifically, this paper seeks to answer the following crucial question: What are the various interesting and non-trivial challenges in PCN research that require immediate attention from the academic and research community?
-By addressing this question, we aim to identify the most pressing problems and future research directions that interested readers can immediately work on.
-Through this analysis, we hope to inspire researchers and practitioners to tackle these challenges to make PCNs more secure and versatile.
+Despite these advantages, the current state-of-the-art in PCNs presents a variety of challenges that require further exploration.
+In this paper, we survey several fundamental aspects of PCNs, such as pathfinding and routing, virtual channels, state channels, payment channel hubs, and rebalancing protocols.
+We aim to provide the reader with a detailed understanding of the various aspects of PCN research, highlighting important advancements.
+Additionally, we highlight the various unresolved challenges in this area.
+Specifically, this paper seeks to answer the following crucial question: What are the various interesting and non-trivial challenges in fundamental infrastructure design leading to efficient transaction processing in PCN research that require immediate attention from the academic and research community?
+By addressing this question, we aim to identify the most pressing problems and future research directions, and we hope to inspire researchers and practitioners to tackle these challenges to make PCNs more secure and versatile.
 Index Terms:
@@
 The cyrptocurrency market is increasing at a rate of $\approx$ 8.00% every year [1].
-This rise in popularity can be attributed to the following reasons: 1) cryptocurrency transactions can be carried without the presence of a trusted entity.
-Fiat currency based transactions on the other hand, require the presence of a trusted financial organization such as a bank.
+This rise in popularity can be attributed to the following reasons: 1) cryptocurrency transactions can be carried out without the presence of a trusted entity.
+Fiat currency transactions on the other hand, require the presence of a trusted financial organization such as a bank.
 2) cryptocurrency transactions do not subject the user to any limits on the number and type of transactions.
-Fiat currency based transactions are limited in their amount and number, depending on several factors such as the currency, geographical location, etc.
+Fiat currency transactions are limited in their amount and number, and depend on several factors such as the currency, geographical location, etc.
 [2].
 Each transaction posted to the Bitcoin blockchain takes around 7 seconds to be validated [3, 4].
-The procedure of validation involves verifying that the transaction posted to the blockchain contains all the required fields and the signature of the user creating the transaction tuple is valid.
+The procedure of validation involves verifying that the transaction posted to the blockchain contains all the required fields and if the signature of the user creating the transaction tuple is valid.
 Once the validation procedure is successfully completed, the transaction is included in a block that would be mined on the blockchain at some time in the future.
@@
 The broadcasting of this revocation to the blockchain prevents the malicious party from stealing funds of the honest party.
-The idea of a payment channel that exists between two parties can be extended to a number of nodes, creating a network of payment channels, called a payment channel networks or PCN.
-The advantage of PCNs is, they enable users that are not connected by a direct payment channel to make payments between each other in an off-chain manner.
+The idea of a payment channel that exists between two parties can be extended to a number of nodes, creating a network of payment channels, called a payment channel network or PCN.
+PCNs enable users that are not connected by a direct payment channel to make payments between each other in an off-chain manner.
 An example PCN is shown in Figure 2.
 In the figure, consider Alice who intends to make a payment to Hector, with whom she does not share a payment channel.
-The naive way to process this transaction would for Alice to open a payment channel with Hector, which involves Alice making an expensive blockchain write for the channel opening.
+The naïve way to process this transaction would be for Alice to open a payment channel with Hector, which involves Alice making an expensive blockchain write for the channel opening.
 Each payment channel opening costs 2.4 USD for blockchain writes [8, 9, 10].
@@
 This difference makes the transaction cost associated with opening a payment channel cheaper than making a payment using a blockchain write., regardless of the transaction amount.
-This cost can be avoided by using the off-chain PCNs, which incurs no additional fees.
+This cost can be avoided by using off-chain PCNs, which incur no additional fees.
 Additionally, transactions on the BTC blockchain take around 2 hours to be confirmed as of June 13, 2024 [5], whereas PCNs can process transactions instantly.
@@
 These figures reflect $\mathsf{LN}$’s size and growth.
+In this SoK, we do not survey the various types of attacks in PCNs [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 21].
+The attacks in PCNs usually have overlaps in their strategy and execution, and most of them currently do not have efficient and fully developed mitigation mechanisms proposed.
+Our conjecture is that, their countermeasures might also have design overlaps as and when they are proposed.
+Hence, we believe attacks in PCNs and their countermeasures require their own taxonomy.
+For this SoK, we have covered papers in various areas of PCNs during the time period of 2019-2024 across Tier-1, Tier-2, Tier-3 security conferences in CS, since the norm for security/privacy research and computer science research in general is peer-reviewed conferences.
 Contributions:
-1) We qualitatively compare the recent work in several aspects of PCNs, viz.
-pathfinding and routing, virtual channels, state channels, payment channel hubs and rebalancing using several relevant properties (metrics) along with providing a reasoning why these metrics have been chosen for comparison.
-2) In this paper, we point out the open problems in all the areas that we survey and we also discuss why solving those problems is a hard research challenge.
+1) We qualitatively compare the recent work in several aspects of building PCNs, viz.
+pathfinding and routing, virtual channels, state channels, payment channel hubs, and rebalancing using several relevant properties (metrics) along with providing a reasoning why these metrics have been chosen for comparison.
+2) We point out the open problems in all the areas that we survey and we also discuss why solving those problems is a hard research challenge.
 Outline: In Section II, we start with describing the concept of pathfinding and routing in PCNs and qualitatively compare work published in that area.
-In Section III, we describe virtual channels which have been proposed to address issues with multi-hop routing in PCNs, and compare work in this area.
-In Section III, we describe and compare state channel protocols, which are a more generalized adaptation of virtual channels and can facilitate the execution of arbitrary applications between nodes in the PCN (not just payments).
+In Section V, we describe virtual channels which have been proposed to address issues with multi-hop routing in PCNs, and compare work in this area.
+In Section IV, we describe and compare state channel protocols, which are a more generalized adaptation of virtual channels and can facilitate the execution of arbitrary applications between nodes in the PCN (not just payments).
 In Section V, we cover payment channel hubs, which are similar to virtual channels and state channels, but facilitate a different use case for payments in PCNs.
 All of the aforementioned PCN mechanisms consume one common resource: the local balance of a node in a PCN.
-In Section IV, we discuss rebalancing, which addresses the important function of replenishing channel funds in the PCN.
-In Section VII, we present the reader with the current research gaps in all of the aforementioned areas and in Section IX, we conclude the paper.
+In Section VI, we discuss rebalancing, which addresses the important function of replenishing channel funds in the PCN.
+In Section VII, we present the reader with the current research gaps in all of the aforementioned areas and also describe why bridging those gaps is hard.
+In Section VIII, we conclude the paper.
 Prior work: Prior works by Khojasteh et al.
@@
 [43] survey the work done only in the area of pathfinding and routing protocols and their privacy aspects in PCNs.
-Whereas, in this paper we cover the entire spectrum of PCN research: rebalancing, virtual channels and state channels, pathfinding and routing.
+Whereas, in this paper we cover the entire spectrum of PCN research: rebalancing, virtual channels, state channels, pathfinding and routing, and tumblers.
 Apart from this [42, 43] do not provide any information about the open problems in PCNs, which we do in our work.
@@
 Furthermore [44] was published in 2020 and does not cover most of the recent work published in PCNs.
-TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs.
-Protocols Concurrency Privacy Topology privacy Avoids source routing Decentralized Atomicity Disjoint graphs Fees
-MPCN-RP [18] $\checkmark$ $\checkmark$ (flat)
-Eckey et al [19] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Vein [20] $\checkmark$ (dynamic)
-Auto tune [21] $\checkmark$ (flat)
-Kadry et al. [22]
-FSTR [23]
-SilentWhispers [24] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Blanc [25] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-SpeedyMurmurs [26] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Spider [27]
-Flash [28] $\checkmark$
-Coinexpress [29] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Webflow [30] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Robustpay [31] $\checkmark$ $\checkmark$ (flat)
-Robustpay+ [32] $\checkmark$ $\checkmark$ (flat)
-RACED [33] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Cheapay [34] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (flat)
-Auroch [35] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (dynamic)
-Yang et al. [36] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-SPRITE [37] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs
+Protocols Concurrency Privacy Topology privacy Avoids source routing Decentralized Atomicity Disjoint graphs Fees Year
+SilentWhispers [45] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2017
+SpeedyMurmurs [46] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2018
+Coinexpress [47] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2018
+Blanc [48] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2019
+Robustpay [49] $\checkmark$ $\checkmark$ (flat) 2019
+Flash [50] $\checkmark$ 2019
+Cheapay [51] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (flat) 2019
+Eckey et al [52] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2020
+FSTR [53] 2020
+Spider [54] 2020
+Vein [55] $\checkmark$ (dynamic) 2021
+Kadry et al. [56] 2021
+Webflow [57] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2021
+Robustpay+ [58] $\checkmark$ $\checkmark$ (flat) 2021
+MPCN-RP [59] $\checkmark$ $\checkmark$ (flat) 2022
+Auto tune [60] $\checkmark$ (flat) 2023
+Yang et al. [61] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2023
+RACED [62] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2024
+Auroch [63] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (dynamic) 2024
+SPRITE [64] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2024
 II Pathfinding and Routing
-Motivation: One of the areas in PCNs which has garnered significant attention from the academic community is pathfinding and routing.
-It is defined as the process of finding a path comprising of several nodes from a sender to a receiver in a PCN along which the payment intended to be sent by the sender can be forwarded to the receiver and forwarding the payment along that path.
-Intuitively, it may seem that well-known pathfinding and routing protocols from the networks domain can be easily applied to PCNs. Unfortunately, there are several problems with this.
-Even though PCNs are decentralized and distributed peer-peer networks, pathfinding and routing protocols of traditional networks cannot be applied to solve the pathfinding and routing problem in PCNs, because of fundamental differences between intent and the network features: 1) Transfer of data and transfer of money: The intent behind traditional network’s pathfinding and routing protocols is to transfer data between connected devices.
-In PCNs, the intent to transfer money between users in a decentralized and distributed manner.
-2) Change in bandwidth and change in monetary state: When data is transferred between connected devices using network pathfinding and routing protocols, the bandwidth, which is represented by the link weights between the nodes does not change.
-However, transfer of money between users in a payment channel permanently alters the state of the sender and receiver involved by subtracting and adding money from and to their link weights respectively.
-3) Latency and routing fees: The cost of sending data in network pathfinding and routing protocols is measured in terms of latency, which is defined as the time delay experienced in transferring data from one device to another.
-The delay in data transfer is experienced by the receiver and the cost of the delay is usually paid by the sender.
-The cost of sending payments in a PCN is measured in terms of money, called as routing fees.
-This encourages greedy behavior and makes the PCN prone to several attack vectors.
-Due to these reasons, pathfinding and routing in PCNs is a fundamentally different and a hard research challenge.
-The properties in Table I represent the fundamental principles of fiat currency transactions and on-chain payment mechanisms, which we want reflected in off-chain payments.
+Motivation: One of the areas in PCNs that has garnered significant attention from the academic community is pathfinding and routing.
+Pathfinding is defined as the process of finding a path comprising several nodes from a sender to a receiver in a PCN along which a payment can potentially be forwarded, and routing is the process of actually forwarding the payment along the found path.
+Intuitively, it may seem that well-known pathfinding and routing protocols from the wired and wireless networks domain can be easily applied to PCNs.
+Unfortunately, there are several problems with this: 1) Traditional networks focus on the transfer of data, PCNs on the other hand, transfer money in a decentralized manner.
+2) Data transfer in traditional peer-peer networks does not alter the bandwidth, whereas money transfer in PCNs alters the monetary state of the nodes involved.
+3) Cost in traditional networks is measured in terms of latency, whereas in PCNs, it involves routing fees, leading to greedy behavior among users and makes PCNs vulnerable to various attacks [44].
+The properties on the columns in Table I represent the fundamental principles of fiat currency transactions and on-chain payment mechanisms, which we want reflected in off-chain payments.
 These properties are generally agreed upon in the literature by several works such as [62, 63, 64, 48, 46, 45] as common evaluation metrics for pathfinding and routing protocols in PCNs.
@@
 In Table I, we present a qualitative comparison of these routing protocols with respect to the properties they achieve.
-Ideal properties:
-1) Concurrency: Concurrency is achieved when a pathfinding and a routing protocol enables the nodes to forward more than one payment simultaneously.
-2) Privacy: Privacy is maintained when a node’s identity is known only to its immediate neighbors and not to the entire network.
+In this paper across all sections, if any prior work has identified a property as ideal or has identified a gap in research, we give an appropriate citation(s).
+If there is no citation provided, it indicates that the corresponding property/research gap has been identified by us.
+Ideal properties: 1) Concurrency: Concurrency is achieved when a pathfinding and a routing protocol enables the nodes to forward more than one payment simultaneously [65].
+Importance: At a given instant of time, many users will be using the PCN to make offchain payments.
+Hence it is important for a routing protocol to support concurrency.
+2) Privacy: Privacy is maintained when a node’s real identity is known only to its immediate neighbors and not to the entire network.
+Importance: Information of a node such as its identity, local balance, connections with other nodes in the network and its transaction history are private and should not be known to anyone else.
 3) Topology privacy: Topology privacy is preserved when no single node has knowledge of the entire network topology.
+Importance: If topology privacy is not preserved, it violates the privacy of every node in the network.
+Making network topology public can potentially lead an adversary to reconstruct transaction paths, which in turn can lead to an adversary selectively targeting a certain set of nodes.
 4) Avoids source routing: Source routing is avoided when the sender does not determine the path to the receiver.
-5) Decentralization: Decentralization is achieved when there are no centralized or trusted entities responsible for constructing paths for senders.
-6) Atomicity: Atomicity is ensured when the payment is routed all the way from the sender to the receiver or the payment is not routed at all.
-7) Disjoint Graph Applicability: A pathfinding and routing protocol is considered applicable to disjoint graphs if it functions even when the network graph consists of islands.
+Importance: If a sender determines the complete path to the receiver, it means that he has access to the entire network topology.
+PCNs are highly dense and dynamic in nature.
+It is practically infeasible for a node to maintain an updated network topology all the time.
+5) Decentralization: Decentralization is achieved when there are no centralized, trusted entities responsible for constructing paths for senders.
+Importance: Cryptocurrency payments made using the blockchain are by nature decentralized, hence routing protocols which facilitate offchain cryptocurrency payments should also be decentralized.
+6) Atomicity: Atomicity is ensured when the payment is routed all the way from the sender to the receiver, or the payment is not routed at all.
+Importance: Atomicity is important since it ensures that honest people do not lose their funds because of malicious behavior by other parties in the system.
+7) Disjoint graphs: A pathfinding and routing protocol is considered applicable to disjoint graphs if it functions even when the network graph consists of islands.
+Importance: PCNs often comprise of islands which only have a couple of nodes.
+These islands are often disconnected from other dense parts of the PCN.
+A routing protocol should be able to facilitate transactions between any pair of nodes irrespective of their location.
 8) Fees: Routing fees is the amount a node charges for forwarding the payment to the next node along a path from the sender to receiver.
@@
 Flat/fixed fees means that the fees charged for routing payments remains the same irrespective of the transaction amount being routed.
-If the fees charged by a node varies according to the transaction amount, it is referred to as dynamic fees, typically a percentage of the amount being transacted.
+If the fees charged by a node varies according to the transaction amount, it is referred to as dynamic fees, typically a percentage of the amount.
+Importance: Every node along a payment path aids the sender in transaction processing by forwarding the payment to the next node along the path.
+The nodes need to be paid in the form of routing fees for their service.
+As illustrated in Table I, routing protocols for PCNs have evolved significantly over the years.
+The two most significant advancements are taking routing fees into consideration and providing support for privacy.
+For instance, $\mathsf{LN}$ provides users with two sets of keys, a long-term keypair and an alias (a temporary identity) helping to conceal their identities and ensure privacy.
+Despite these developments, two overarching research problems remain that require attention.
+We discuss them in detail in Section VII.
 III Virtual channels
 (a) Virtual channel example [66, 67]
-(b) Multi-hop virtual channel (the dots between Jill and Alice depict several intermediate nodes)
-(c) Recursive virtual channel
+(b) Recursive virtual channel
+(c) Multi-hop virtual channel (the dots between Jill and Alice depict several intermediate nodes)
 Figure 3: Different types of virtual channels
 TABLE II: Comparison of Virtual Channel Protocols
-Protocol BC Validity Fee Privacy Off-chain dispute resolution Recursive Multihop
-Perun [38] TC Limited
-Jourenko et al. [40] UTXO Limited $\checkmark$
-Aumayr et al. [41] UTXO Limited $\checkmark$ (fixed)
-Elmo [42] UTXO Unlimited $\checkmark$
-Donner [43] UTXO Unlimited $\checkmark$ (time based) $\checkmark$ $\checkmark$
-Jia et al. [44] UTXO and TC Limited $\checkmark$ (fixed)
-Eckey et al. [45] TC Limited $\checkmark$
-Generalized state channels [46] TC Limited $\checkmark$
-Transactions in PCNs are routed from the sender to the receiver using a path of intermediate nodes.
-This pathfinding and routing mechanism requires the nodes along the payment path to be available for a transaction to be processed.
-Nodes, however, sometimes can choose to go offline or there can be network/service disconnections.
-This might cause the transaction to fail.
-Virtual channels, which are built on top of existing payment channels, provide a solution to this problem.
-Virtual channels facilitate payments between a pair of nodes using an intermediate node.
-The intermediate node has individual payment channels open with the other two nodes.
-The intermediary and the pair of nodes lock coins with each other in their respective payment channels to establish a virtual channel.
-Once the virtual channel has been established, the two nodes can be involved in a unlimited number of payments.
-These payments can be processed even if the intermediate node goes offline.
-The first virtual channel construction, Perun, was proposed by Dziembowski et al.
-in [38].
-Assume there are two users, Alice and Bob, along with an intermediary as depicted in Figure 3(a).
-Alice establishes a payment channel, $\alpha_{A}$, with the intermediary by locking $Y_{A}$ coins as collateral; the intermediary contributes $Y_{I}$ coins into this channel.
-Similarly, Bob locks $Z_{B}$ coins into his channel, $\alpha_{B}$, with the intermediary, which also deposits $Z_{I}$ coins from its possession.
-This marks the completion of the opening phase, effectively locking the coins belonging to Alice and Bob.
-Following this, Alice and Bob create a virtual channel to facilitate transactions between them.
-The virtual channel will include coins deducted from both Alice’s and Bob’s payment channels.
-Specifically, $X_{A}$ coins will be subtracted from $Y_{A}$ on Alice’s side and deposited into the virtual channel, while $X_{B}$ coins will be subtracted from $Z_{B}$ and will be deposited from Bob’s side.
-Now Alice and Bob can be involved in an unlimited number of payments without the intermediary being online.
-The idea of a virtual channel between a pair of nodes involving a single intermediary/intermediate node has been extended to establishing a virtual channel recursively over several hops involving several intermediaries, leading to the construction of a recursive virtual channel [40, 42, 43].
-An example of a recursive channel is shown in the Figure 3(c).
-In the figure Trent and Alice are a pair of nodes who intend to establish a virtual channel between them.
-They do in three steps as follows; 1) Trent and Hu construct a virtual channel $V_{1}$ with Rajiv as an intermediary.
-2) Hu and Alice construct a virtual channel $V_{2}$ with Jill as an intermediary.
-3) Trent and Alice establish a virtual channel between them $V_{3}$.
+Protocol BC Validity Fee Privacy Off-chain dispute resolution Recursive Multihop Year
+Generalized state channels [68] TC Limited $\checkmark$ 2018
+Eckey et al. [69] TC Limited $\checkmark$ 2019
+Perun [66] TC Limited $\checkmark$ 2019
+Jourenko et al. [70] UTXO Limited 2020
+Aumayr et al. [71] UTXO Limited $\checkmark$ (fixed) 2021
+Elmo [72] UTXO Unlimited $\checkmark$ 2021
+Donner [73] UTXO Unlimited $\checkmark$ (time based) $\checkmark$ $\checkmark$ 2023
+Jia et al. [74] UTXO, TC Limited $\checkmark$ (fixed) 2023
+Motivation: Transactions in PCNs are routed from the sender to the receiver using a path of intermediate nodes.
+Current pathfinding and routing mechanisms require the nodes along the payment path to be available for a transaction to be processed.
+However, nodes can sometimes choose to go offline or there can be network/service disconnections causing transaction failures.
+Furthermore, each node along a payment path charges its own fees for forwarding the payment, which is paid by the sender and increases with the path length, hence the time taken to route a payment grows linearly in the path length.
+Virtual channels, which are built on top of existing payment channels, solve these problems.
+Initial constructions of virtual channels facilitated payments between a pair of nodes using a single intermediate node [66].
+The intermediate node needs to have individual payment channels open with the other two nodes.
+The intermediary and the pair of nodes lock coins with each other in their respective payment channels and a virtual channel is established.
+Upon establishment of the virtual channel, the pair of nodes can be involved in a unlimited number of payments.
+These payments can be processed without the intermediate node being online.
+It might be better to use virtual channels from a routing fees perspective, since unlike routing protocols, nodes in virtual channels do not charge a routing fee for every transaction.
+Alice, Bob, and an intermediary establish a virtual channel as shown in Figure 3(a).
+Alice locks $Y_{A}$ coins and the intermediary locks $Y_{I}$ coins in the payment channel $\alpha_{A}$.
+Similarly, Bob locks $Z_{B}$ coins and the intermediary locks $Z_{I}$ coins in their channel $\alpha_{B}$.
+The virtual channel $V$ is created once Alice locks $X_{A}$ coins from her balance in $\alpha_{A}$ and Bob locks $X_{B}$ coins from his balance in $\alpha_{B}$.
+Now Alice and Bob can process payments without the intermediary’s online presence.
+The idea of a virtual channel between a pair of nodes involving a single intermediary has been extended to establishing a virtual channel recursively over several hops involving several intermediaries, leading to the construction of a recursive virtual channel [70, 72, 73].
+An example of a recursive channel is shown in Figure 3(b).
+A recursive virtual channel enables transaction processing between any pair of nodes, by recursively establishing virtual channels between several intermediaries, as opposed to [66], which facilitates virtual channels only between a pair of nodes connected directly to an intermediary.
 A multihop virtual channel is constructed by establishing a single virtual channel between a pair of users over a path comprising of several intermediate nodes.
-An example of multihop virtual channel is shown in the Figure 3(b).
-Rajiv and Alice establish a single virtual channel $V$ between them.
-Motivation: We have already described how PCNs offer a solution to the blockchain scalability problem by facilitating instantaneous payments in an off-chain manner.
-PCNs however, have few limitations: 1) The path generated in a PCN to process a payment between a sender and a receiver usually contains one or several intermediate nodes.
-All the nodes along the path are required to be available to route the payment until the receiver.
-This may not be possible at all times since, nodes can be disconnected from the network due to various reasons such as power outages, DoS attacks, etc.
-In a virtual channel, once the coins of the intermediary(ies) are locked, it can choose to remain offline without affecting the transaction processing.
-2) Each node along the path from the sender to the receiver charges a fees called as routing fees/forwarding fees for forwarding the payment to the next node along the path.
-The sender has to pay all the nodes their respective forwarding fee along the path.
-This increases the overall cost of routing the payment.
-Virtual channels provide a solution to this problem, since they do not charge fees on a per transaction basis.
-Instead, the fees is charged based on a time epoch basis (the time for which the coins of the intermediary/intermediaries are locked in the virtual channel) or a fixed fee can be agreed in advance by all the parties involved in the virtual channel construction.
-3) The time taken to route a payment from a sender to a receiver is directly proportional to the path length.
-A longer path would result in an increase in the overall pathfinding and routing time.
-Virtual channels do not process payments through a path comprising of several intermediate nodes, hence the overall transactions’ pathfinding and routing time does not depend on the number of intermediate nodes between the sender and the receiver.
+An example of multihop virtual channel is shown in Figure 3(c).
+A multihop virtual channel is an improvement over recursive virtual channels.
+In a multihop virtual channel, a single virtual channel can be established between any pair of nodes which are separated by several intermediaries.
 Virtual channels should not add an unnecessary burden to users, and should mirror the operations of payment channels as closely as possible, with the added benefit of no on-chain transactions at all, while maintaining comparable security/privacy properties.
-We now give the properties desired from an ideal virtual channel, and compare the works in this area in Table II on the extent to which they achieve these properties.
-Ideal properties:
-1) BC: BC is the blockchain on which the corresponding virtual channel protocol can be deployed.
-TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO based scripting mechanism such as Bitcoin.
-Ideally, a virtual channel protocol should be deployable on any blockchain without the need for any additional programming/scripting requirements.
-The virtual channel protocol should not require any modifications to the existing consensus mechanisms of the underlying blockchains.
-2) Validity: This determines the validity of the virutal channel.
-A limited validity means that the virtual channel is valid for a predetermined time period (which is decided by parties involved in the virtual channel).
-Unlimited validity means that the virtual channel can stay open until the parties involved decide to initiate closing.
-This provides better and efficient transaction processing between users who intend to have frequent payments made between each other.
-3) Fee: This metric determines if the virtual channel takes into account the fees charged by the intermediate node(s) involved in the channel’s establishment.
-The fixed fee model implies that a predetermined, fixed fee is paid to the intermediate node(s) which is agreed upon by all the parties in the virtual channel and the time based fee model implies that a fixed fee is paid to the intermediary which depends on the time for which the virtual channel stays open.
-Having an well-defined fee structure will motivate the intermediary/intermediaries to participate in virtual channel creation.
-4) Privacy: Similar to payment channels, we want privacy of nodes for virtual channels, which is said to be achieved when the identities of the nodes in the virtual channel protocols are not known to anyone except for their immediate neighbors.
-5) Offchain dispute resolution: This metric determines if the transaction disputes in a virtual channel require a blockchain write.
-6) Recursive: A virtual channel is said to be recursive or hierarchical if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing virtual channels on top of existing virtual channels from sender to receiver.
-We give an example of a recursive virtual channel in Figure 3(c).
-7) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel between any pair of intermediate nodes along the path.
-We give an example of a multihop virtual channel in Figure 3(b).
+We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
+Blockchain terminology: In the rest of the paper BC denotes a blockchain.
+TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO-based scripting mechanism such as Bitcoin.
+In Table II, we present a qualitative comparison of virtual channel protocols.
+Ideal properties: 1) Validity: This determines the validity of the virtual channel.
+A limited validity means that the virtual channel is valid for a predetermined time period (which is decided by nodes involved in the virtual channel).
+Unlimited validity means that the virtual channel can stay open until the nodes involved decide to initiate closing [69].
+Importance: Unlimited validity provides better and efficient transaction processing between users who intend to have frequent payments made between each other.
+2) Fee: This metric determines if the virtual channel takes into account the fees charged by the intermediate node(s) involved in the channel’s establishment.
+The fixed fee model implies that a predetermined, fixed fee is paid to the intermediate node(s) which is agreed upon by all the nodes in the virtual channel before the channel establishment.
+The time-based fee model implies that the fee paid to the intermediary depends on the time for which the virtual channel stays open [69].
+Importance: Having an well-defined fee structure will motivate the intermediary/intermediaries to participate in virtual channel creation.
+3) Privacy: In any virtual channel construction, the real identity of a node should only be known to its immediate neighbor(s) [66].
+Importance: Privacy is important since it helps in preserving topology privacy.
+4) Offchain dispute resolution: This metric determines if the transaction disputes in a virtual channel require a blockchain write.
+Importance: Ideally, we would want a virtual channel construction to have off-chain dispute resolution since blockchain writes are expensive and time-consuming.
+5) Support for multihop virtual channels with several intermediaries: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channels between any pair of intermediate nodes along the path.
+Importance: This property is ideal since it facilitates payments between any pair of nodes in the network by establishing only a single virtual channel, as opposed to a recursive virtual channel in which a sender/receiver, and the intermediate nodes lock coins in multiple virtual channels at the same time.
+The most significant developments in virtual channels over the years are that newer protocols have incorporated a fees to be paid for the intermediary(ies) that lock coins in virtual channels and virtual channels now offer support for both TC based and UTXO based blockchains.
+Despite these developments, efficient virtual channel protocol design has three overarching research problems, which are discussed in Section VII.
 (a) State channel opening
@@
 TABLE III: Comparison of State Channel Protocols
-Protocol BC General/purpose state channels Security Graceful exit Off-chain dispute resolution
-Pisa [47] TC $\checkmark$
-Hydra [48] UTXO $\checkmark$ $\checkmark$ $\checkmark$
-Sprites [49] UTXO $\checkmark$ $\checkmark$ $\checkmark$
-Aumayr et al. [50] UTXO, TC $\checkmark$ $\checkmark$
-Origami [51] TC $\checkmark$ $\checkmark$
-Forcemove [52] TC $\checkmark$
+Protocol BC General/purpose state channels Security Graceful exit Off-chain dispute resolution Year
+Forcemove [75] TC $\checkmark$ 2018
+Pisa [76] TC $\checkmark$ 2019
+Sprites [77] UTXO $\checkmark$ $\checkmark$ 2019
+Hydra [78] UTXO $\checkmark$ $\checkmark$ 2021
+Aumayr et al. [79] UTXO, TC $\checkmark$ 2021
+Origami [80] TC $\checkmark$ $\checkmark$ 2023
 IV State channels
@@
 This leads to the PCN becoming more versatile.
-State channels should be deployable on any blockchain regardless of the underlying programming/scripting requirements and should be able to facilitate the execution of any decentralized application between the parties involved, while maintaining comparable security properties (in comparison to payment channels) and should ideally be able to resolve disputes without accessing the blockchain.
+State channels should be deployable on any blockchain regardless of the underlying programming/scripting requirements and should be able to facilitate the execution of any decentralized application between the parties involved, while maintaining comparable security properties (to payment channels) and should ideally be able to resolve disputes without accessing the blockchain.
 We now give the properties desired from an ideal state channel, and compare the works in this area on the extent to which they achieve these properties.
@@
 We describe the metrics used for comparison below.
-Ideal properties:
-1) Blockchain: This is the blockchain on which the corresponding state channel protocol can be deployed.
-TC is a blockchain that supports a Turing complete programming language and UTXO if a blockchain that supports a UTXO based scripting.
-2) General purpose state channels: This is the ability of the state channel to facilitate the execution of any application in an off-chain manner by using a single application-agnostic instance of state channel on the blockchain.
-3) Security: Security is achieved by a state channel protocol if no honest party loses funds because of malicious behavior by other parties in the protocol.
-4) Graceful exit: A state channel protocol is said to employ graceful exit if it has clear and well-defined mechanisms for nodes joining or leaving the state channel.
-5) Off-chain dispute resolution: This metric indicates if a dispute that occurs between the parties involved in the state channel can be resolved without accessing the blockchain.
+Ideal properties: 1) General purpose state channels: This is the ability of the state channel to facilitate the execution of any application supported by the underlying blockchain in an off-chain manner Importance: This is an ideal feature since it facilitates the deployment of any application without accessing the blockchain, making state channels more versatile.
+2) Graceful exit: A state channel protocol is said to employ a graceful exit if it has clear and well-defined mechanisms for nodes joining or leaving the state channel without accessing the blockchain.
+Importance: This is an important property since most of the state channel protocols require the user to interact with the blockchain when they join/leave the state channel(s) they are a part of.
+3) Off-chain dispute resolution: This property indicates if a dispute that occurs between the parties involved in the state channel can be resolved without accessing the blockchain.
+Importance: Ideally, we would want a state channel construction to have off-chain dispute resolution since blockchain writes are expensive and time-consuming.
 V Tumblers
-TABLE IV: Comparison of Tumblers
+TABLE IV: Comparison of Payment Channel Hubs
 Protocol BC
@@
 aborts
-Time epochs
+Independent of
+epochs
 Dynamic
@@
 amount
-BOLT [53] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Nocust [54] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-MIXCT [55] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-A2L [56]
-UTXO and
-TC
-$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Blindhub [57]
-UTXO and
-TC
-$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Teechain [58] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Accio [59] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Turbo [60] TC $\checkmark$ $\checkmark$
-Boros [61] TC $\checkmark$ $\checkmark$ $\checkmark$
-Tumblebit [62] UTXO $\checkmark$ $\checkmark$ $\checkmark$
-Motivation: A payment channel hub is a layer-2 mechanism that facilitates transaction processing between a sender/receiver pair that do not have a payment channel open between them.
+Year
+BOLT [81] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2017
+Tumblebit [82] UTXO $\checkmark$ $\checkmark$ 2017
+Nocust [83] TC $\checkmark$ $\checkmark$ $\checkmark$ 2018
+Teechain [84] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2019
+$A^{2}L$ [85] UTXO, TC $\checkmark$ $\checkmark$ $\checkmark$ 2021
+Accio [86] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2021
+Boros [87] TC $\checkmark$ $\checkmark$ $\checkmark$ 2021
+MIXCT [88] TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2022
+Turbo [89] TC $\checkmark$ 2022
+Blindhub [90] UTXO, TC $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2023
+Motivation: A payment channel hub (tumbler) is a multi-party off-chain system where users can establish payment channels with a central hub, which acts as an intermediary.
+It allows multiple users to send payments to each other without the need for direct payment channels between each user pair.
+The hub coordinates payments between different participants.
 The intermediary which facilitates payments is called a tumbler.
-The tumbler has multiple payment channels open with several nodes in the network to facilitate payments.
-The hubs can potentially be useful in facilitating micropayments between a buyer and seller, without either of them needing to open a payment channel with each other.
-Payment channel hubs can be classified into two types: on-chain and off-chain.
-Early hubs, also called hubs or mixers, were on-chain [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80], but they all suffered from scalability issues due to having to post each transaction on the blockchain.
-This led to the development of offchain payment channel hubs for specific blockchains, e.g., Bolt [53] for Zcash, Nocust [54] and MixCT [55] for Ethereum.
+Though a payment channel hub uses the same underlying infrastructure as that of PCNs and virtual channels, each of these constructions have their own use cases.
+PCNs are usually used when two nodes Alice and Bob transact on an infrequent basis.
+Virtual channels are used if Alice and Bob transact frequently, e.g., if Bob provides Alice with a service every month.
+Payment channel hubs are used when Alice needs to pay several receivers on a frequent basis and she does not want the tumbler to know the receivers.
+Payment channel hubs can be classified into two types: onchain and off-chain.
+Early hubs, also called hubs or mixers, were on-chain [91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102], but they all suffered from scalability issues due to having to post each transaction on the blockchain.
+The scalability issues of on-chain payment channel hubs have led to the development of offchain payment channel hubs for specific blockchains, e.g., Bolt [81] for Zcash, Nocust [83] and MixCT [88] for Ethereum.
 The most general-purpose payment channel hubs are Tumblebit [82], $A^{2}L$ [85], and Blindhub [90].
@@
 We now give the properties desired from an ideal payment channel hub, and compare the works in this area on the extent to which they achieve these properties.
-We give a qualitative comparison of off-chain payment channel hubs in the Table IV.
-We describe the comparison metrics below:
-Ideal properties:
-1) Blockchain: BC is the blockchain on which the corresponding Tumbler protocol can be deployed.
-TC is a blockchain that supports a Turing complete programming language, such as Ethereum, and UTXO is a blockchain that supports the UTXO based scripting mechanism such as Bitcoin.
-2) Relationship anonymity: It ensures that the relationship between a sender and a receiver for a given transaction should not be known to the tumbler.
-3) Privacy against aborts: Tumbler should not be able to deduce the identities of a sender/receiver in case of a transaction abort, regardless of which party is responsible for the abort.
-4) Time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch.
-A payment channel hub that supports transaction processing based on time epochs is not efficient because nodes cannot process payments between time epochs, which can be of arbitrary length.
+In Table IV, we qualitatively compare several off-chain tumblers.
+Ideal properties: 1) Relationship anonymity: It ensures that the relationship between a sender and a receiver for a given transaction should not be known to the tumbler [81].
+Importance: This is an important property since one of the main uses and design goals of a tumbler is to facilitate anonymous transactions.
+Hence, any tumbler construction should satisfy this property.
+2) Privacy against aborts: Tumbler should not be able to deduce the identities of a sender/receiver in case of a transaction abort, regardless of which party is responsible for the abort [90].
+Importance: This is an important property to have since aborts can happen due to several reasons such as network disconnections, power outages, etc.
+Malicious nodes can also deliberately abort a protocol.
+In the case of tumbler protocols, the nodes whose payments did not succeed can be linked to each other once the protocol execution completes.
+3) Independent of time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch [81].
+This is not ideal.
+Importance: A payment channel hub that supports transaction processing based on time epochs is not efficient because nodes cannot process payments between time epochs, which can be of arbitrary length.
 The duration between time epochs depends on several factors such as the unit of the time epoch (absolute time or relative time measured in terms of block height), number of users that will most likely involved in the next time epoch, etc.
-5) Dynamic corruption: The payment channel hub should be able to handle the deviation of any party from the protocol at any point during its execution.
-i,e the set of adversaries should not be static during the protocol execution.
-6) Atomicity: Atomicity ensured when a payment is routed all the way from the sender to the receiver or the payment is routed at all.
-7) Value privacy: Value privacy is ensured by a payment channel hub, when, given a transaction amount, the tumbler cannot link it to a sender/receiver pair.
-8) Variable amount: The tumbler should be able to process transactions of any value.
-TABLE V: Comparison of Rebalancing Protocols
-Protocol BC Trusted entity Graph compatability Privacy
-Revive [81] TC $\checkmark$ (C) Cycles only
-Hide & Seek [82] UTXO $\checkmark$ (D) Cycles only $\checkmark$
-Subramanian et al. [83] TC and UTXO Agnostic $\checkmark$
-Cycle [84] TC Cycles only $\checkmark$
-Rebal [85] UTXO Cycles only $\checkmark$
-Shaduf [86] TC Agnostic $\checkmark$
-Chen [87] et al. UTXO Cycles
-Musketeer [88] UTXO Cycles
+4) Protection against dynamic corruption of nodes: Current tumbler protocols corrupt nodes at the beginning of protocol execution and assume that the set of corrupted nodes remains constant until the protocol execution terminates.
+Ideally, a payment channel hub should be able to handle the deviation of any party from the protocol at any point during its execution.
+Importance: This is an important property since its infeasible from a practical standpoint to assume that the set of adversaries in the payment channel hub protocol remains constant during execution.
+5) Atomicity: Atomicity is ensured if either the payment is routed all the way from the sender to the receiver or the payment is routed at all.
+Importance: This is important since it prevents honest parties from losing their funds because of malicious parties.
+6) Value privacy: Value privacy is ensured by a payment channel hub, when, given a transaction amount, the tumbler cannot link it to a sender/receiver pair [81].
+Importance: Value privacy is important since only the sender and receiver should know the amount being transacted.
+In the event of every sender/receiver pair transacting a unique amount, lack of value privacy will enable the tumbler to link a sender/receiver pair to a particular transaction, violating relationship anonymity.
+7) Variable amount: The tumbler should be able to process transactions of any amount [81].
+Importance: A tumbler having the ability to process amounts of variable value is important since its practically infeasible to assume that all sender/receiver pairs will always transact a fixed amount.
+Over the years, various payment channel hub constructions have been developed to address specific challenges based on their unique design goals.
+However, we have identified three overarching research problems in payment channel hubs that require attention.
+We discuss them in Section VII.
+TABLE V: Comparison of Rebalancing Protocols; C is centralized and D is distributed
+Protocol BC Trusted entity Graph compatability Privacy Year
+Revive [103] TC $\checkmark$ (C) Cycles only 2017
+Subramanian et al. [104] TC and UTXO Agnostic $\checkmark$ 2019
+Rebal [105] UTXO Cycles only $\checkmark$ 2021
+Hide & Seek [106] UTXO $\checkmark$ (D) Cycles only $\checkmark$ 2022
+Cycle [107] TC Cycles only $\checkmark$ 2022
+Shaduf [108] TC Agnostic $\checkmark$ 2022
+Musketeer [109] UTXO Cycles 2023
+Chen [110] et al. UTXO Cycles 2024
 VI Rebalancing
-Motivation:Rebalancing in payment channel networks involves redistributing the balance within an existing payment channel or across multiple channels to maintain or improve the network’s liquidity and usability.
+Motivation: Rebalancing in PCNs involves redistributing the balance within an existing payment channel or across multiple channels to maintain or improve the network’s liquidity and usability.
 This is done to ensure that both parties in the channel have sufficient funds on their respective sides to send and receive payments.
@@
 A rebalancing protocol should be be able to replenish the link weight of dormant nodes in the PCN irrespective of the PCNs topology, while not having to employ any trusted entity and should not compromise privacy of nodes.
-We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
+We now give the properties desired from an ideal rebalancing protocol, and compare the works in this area on the extent to which they achieve these properties.
 We give a qualitative comparison of rebalancing protocols in Table V.
 We describe the metrics as follows:
-Ideal properties:
-1) Blockchain: This is the blockchain on which the corresponding rebalancing protocol can be deployed.
-As it can be inferred from the Table, the protocols can either be deployed on the Ethereum or the Bitcoin blockchains.
-2) Trusted entity: This metric determines the nature of trust required among parties to deploy the rebalancing protocol.
-A centralized trust (C) indicates the presence of a single trusted entity and distributed trust (D) indicates the distribution of trust across various entities.
-3) Graph compatibility: This indicates the nature of the topology of the network on which a rebalancing protocol can be deployed.
-Cycles indicate that the rebalancing protocol can only deployed on cyclic graph and agnostic indicates that the rebalancing protocol can be deployed irrespective of the network topology.
-4) Privacy: Privacy is said to be achieved when sensitive information such as the local balance and the identities of nodes are not known to anyone except for their immediate neighbors.
+Ideal properties: 1) Trusted entity: This metric determines the nature of trust required among parties to deploy the rebalancing protocol.
+Centralized trust (C) indicates the presence of a single trusted entity and distributed trust (D) indicates the distribution of trust across various entities.
+Importance: It is ideal to have distributed trust for a rebalancing protocol since, PCN payments are by nature decentralized.
+2) Graph compatibility: This indicates the nature of the topology of the network on which a rebalancing protocol can be deployed.
+Cycles indicate that the rebalancing protocol can only deployed on cyclic graphs and agnostic indicates that the rebalancing protocol can be deployed irrespective of the network topology.
+Importance: Ideally a rebalancing protocol should be deployable irrespective of the topology of the graph since its practically infeasible to assume that every node in a PCN will be a part of a cycle.
+3) Privacy: Privacy is said to be achieved when sensitive information such as the local balance and the identities of nodes are not known to anyone except for their immediate neighbors.
+Importance: The importance of privacy for layer-2 protocols has already been described in Section II.
+Rebalancing is a relatively mature area and significant challenges have already been addressed.
 VII Research Gaps & Open Problems
-In this section, we highlight the gaps in research published until now in the areas of pathfinding and routing, virtual channel construction, payment channel hubs, and state channels.
-This section provides an overview of the current research problems in these areas along with a discussion on why is developing a solution to these problems an interesting research challenge.
-VII-A Gaps in pathfinding and routing for PCNs
-We now identify the gaps in research in pathfinding and routing protocols for payment channel networks.
-1) Super node liquidity validation in PCNs: A super node is a highly connected node (it has payment channels open with numerous other nodes in the network) and also has a high liquidity in those channels.
-These super nodes aid other nodes in pathfinding and routing of payments.
-These super nodes are called as a trampoline nodes in $\mathsf{LN}$ [89, 90, 91, 92].
-Several other works in the literature refer to super nodes as routing node, routing helper, landmark node, routers, etc [25, 24, 27, 37, 33].
-One of the main problems with the super nodes is that the sender has no way to know whether the super node possess the local balance (liquidity) to route the payment of the sender.
-The local balance of a super node or any node in a PCN is a private value and should not be known to any node except for its immediate neighbor.
+In this section, we highlight the gaps in research published up until now in the areas of pathfinding and routing, virtual channel construction, state channels and payment channel hubs.
+The gaps are described as research questions, denoted by RQ.
+RQ1: Why is super node liquidity validation in PCNs hard?
+A super node, variously called as a trampoline node, routing node, routing helper, landmark node, router, etc.
+[48, 45, 54, 64, 62] is a highly connected node with numerous high liquidity channels, that helps in pathfinding and routing payments.
+One of the main problems with the super nodes is that a sender has no way to know whether the super node possesses enough liquidity on its channels to route a payment.
+The local balance of a super node in a given channel (or of any node in a PCN) is a private value and should not be known to any node except for its immediate neighbor that it shares the channel with.
 Currently, if a super node does not have enough liquidity to route the payment of a sender, the transaction fails and it has to be retried by the sender.
@@
 The main goal of PCNs is to facilitate instantaneous payments and these transaction retries render such payments almost impossible.
-It might greatly benefit the sender if the it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns in the network.
-Achieving this is hard because the local balance a node has with its immediate neighbors in a PCN is considered private information and the sender has to know whether the super node possess enough liquidity without comprising privacy.
-2) Channel verification in a PCN: To be a part of any PCN, nodes will open payment channels with other nodes in the network.
+It will greatly benefit the sender if it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns.
+RQ2: Why is channel verification in a PCN hard?
+To be a part of any PCN, nodes will open payment channels with other nodes in the network.
 Two nodes open a payment channel between them by posting a transaction to the blockchain.
-This transaction can be posted on the blockchain (as in UTXO based blockchains like BTC) or this transaction can be posted as a function call to an existing smart contract (as in blockchains that support smart contracts like ETH).
-In the most popular PCN, $\mathsf{LN}$, the channel verification procedure is very inefficient.
-It involves the node who intends to verify a channel to request the block in which the channel opening transaction has been included, verifying whether the transaction has been successfully executed by the validator/miner and finally verifying if the channel opening transaction corresponds to a 2-2 multi signature address on the blockchain.
+This transaction can be posted on the blockchain or as a function call to an existing smart contract.
+In the most popular PCN, $\mathsf{LN}$, the procedure of verifying whether a payment channel really exists on the blockchain is very inefficient.
+A node who wants to verify a channel needs to request the block in which the channel opening transaction has been included, verifying whether the transaction has been successfully executed by the validator/miner and finally verifying if the channel opening transaction corresponds to a 2-2 multi signature address on the blockchain.
 The verifier performing these steps is inefficient since all these steps will have already been performed by the miner.
-According to the current verification process, any node would need to perform all of these steps for each payment channel it has open in the network.
-Consider the following scenario: a node $j$ in the PCN has payment channels open with $m$ nodes in the PCN.
-Furthermore, assume that each neighbor opens $k$ new channels for each time epoch.
-For this example, assume the value of the time epoch is 1 block.
-This would mean that the total number of channels the node $j$ has to verify a total of $m*k$ channels for every block.
-This requires the node $j$ to access the blockchain a total of $m*k$ times to perform channel verification.
 Finding a way to do this without blockchain access and in a blockchain agnostic manner in a hard research challenge.
-VII-B Gaps in Tumblers
-We will now describe the research gaps in tumblers.
-1) Privacy against aborts: In a given a transaction, the tumbler should not be able to correctly identify the sender/receiver pair involved.
-Currently, there is only one construction [59] that guarantees this.
-However, this construction works only for blockchains that support Turing complete programming languages.
-Currently, there is no such construction for UTXO based blockchains that guarantees privacy against aborts.
-This is hard because if the sender aborts a transaction, the tumbler will know which receiver did not receive the payment.
-This is hard because, a tumbler mixes multiple transactions together to ensure unlikability among payments.
-Also, a tumbler should enforce atomicity so that the payment either goes through completely or does not go through at all.
-In the case of an abort, an adversary who controls the tumbler might be able to deduce which receiver amongst the ones involved did not receive a payment and will be able to link the corresponding transaction amount to that receiver.
-2) Time epochs: A tumbler should be able to facilitate payments without having a predefined time interval or epoch.
-Currently, the majority of payment channel hub constructions process payments in discrete epochs to increase the size of their anonymity sets.
-An anonymity set refers to a set of users (senders/receivers) whose identity is indistinguishable from each other.
-The size of this set is proportional to the privacy of sender/receiver in the transaction.
-When the interval between two consecutive time epochs is sufficiently large, it will promote more users to take part in the transactions within a time epoch, which improves privacy.
-Not having time epochs results in smaller anonymity sets and hence could possibly weaken privacy.
-The trade off is between privacy and flexibility.
-3) Dynamic corruption: A tumbler should ideally be resistant to dynamic corruption.
-This means the protocol should be resilient to parties deviating from their ideal execution during the beginning, progress, and ending of the protocol’s execution.
-Current tumbler constructions only provide support for static corruption, which may be unrealistic from a practical standpoint.
-Security against dynamic corruption is hard to achieve because the tumbler mixes several transactions together to preserve anonymity.
-Since the tumbler should not be able to link transaction amount to a sender/receiver pair (ideally), if the sender/receiver becomes malicious during the protocol’s execution, ideally, that transaction must be aborted to prevent financial loss to the honest party.
-However, the tumbler should not be able to link a transaction to a sender/receiver pair, making security against dynamic corruption a hard research problem.
-VII-C Gaps in Virtual Channels
-We will now describe the open problems in the area of virtual channels.
-1) Blockchain agnostic: A virtual channel should ideally be blockchain agnostic, meaning it should be deployable regardless of the underlying scripting or programming language structure supported by the blockchain.
-2) Well defined fee structure: The intermediary/intermediaries involved in the virtual channel construction additionally lock coins apart from the ones locked in the underlying payment channel.
-It is necessary to have a well-defined fee structure for this additional service.
-Currently, nodes get paid a routing fees/forwarding fees for every transaction they process.
-This fee is dependent on the transaction amount being processed.
-In the case of virtual channels, having a well-defined fee structure is difficult due to the following reasons: 1) The fee structure should take into account the amount of funds and the time for which these funds of the intermediary/intermediaries are locked in virtual channel.
-2) It also needs to take into account the routing/forwarding fees an intermediary could have earned by not locking up the coins in the virtual channel.
-The fees for an intermediary should be ideally be more than the routing fees it would have earned in order to motivate nodes to be involved in virtual channel construction.
-Currently, no virtual channel construction has a well-defined fee structure.
-3) Off-chain dispute resolution: Ideally a virtual channel protocol should provide the flexibility to resolve any disputes without accessing the blockchain.
-We have seen in Section I, that blockchain writes are expensive.
-Having an off-chain dispute resolution is a hard research challenge due to the following reasons; 1) Any form of layer-2 disputes are usually resolved on the blockchain.
-Since underlying consensus in blockchain usually guarantees resolution of a dispute raised within a specific time period.
-In an off-chain scenario, there is no such mechanism which guarantees this.
-None of the virtual channel constructions have a provision of off-chain dispute resolution.
-4) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel(s) between any pair of intermediate nodes along the path.
-Currently, no VC construction has this capability.
-This is hard because, multihop virtual channel involves designing a mechanism for sequentially locking coins across a path comprising of several intermediate nodes.
-Furthermore, the construction of a multihop virtual channel should also ensure that any transaction does not effect the overall balance neutrality (no node apart from the receiver along the path gains or loses any money).
-Also, the goals of privacy, atomicity, offchain dispute resolution need to be enforced.
-VII-D Gaps in state channel protocols
-We will now describe the gaps in the area of state channel protocols.
-1) General purpose state channels: An ideal blockchain should be able to facilitate deployment and execution of several decentralized applications without requiring the users to access the blockchain for any purpose other than opening and closing the state channel.
-Currently, there are two such constructions [49, 48] which incorporate code reusability into their design.
-These two constructions support UTXO based blockchains and have no provision for offchain dispute resolution.
-2) Offchain dispute resolution: Current state channel protocols use an onchain transaction or a function call to the onchain smart contract in the case of dispute resolution.
+RQ3: Why is designing pathfinding protocols for PCNs, that comprise of several distinct well-connected components a hard problem?
+Though solutions such as [62] exist that solve this problem to a certain extent by using routing helpers/trampoline nodes, the aspect that makes it hard is to quantify the denseness/sparseness of a well-connected component.
+If we assume that we deploy one trampoline node for each well-connected component, there may be well-connected components in the PCN which only have a few nodes (i.e., islands).
+Selection of a trampoline node amongst these nodes is difficult.
+RQ4: Why is designing a routing protocol that supports concurrent payments and is resilient to channel gaming a hard problem?
+Processing concurrent transactions requires the design of a mechanism that allows a node to lock a portion of its liquidity in a channel with an immediate neighbor for one transaction while simultaneously using the remaining liquidity to process another.
+Though there are protocols that support concurrency [47, 45, 48, 62, 63, 64], they are not resilient to the presence of potentially malicious nodes in the PCN, which may initiate transactions with the sole intent of locking liquidity, leading to congestion and disruption in the network.
+RQ5: Why is having a well-defined fee structure for virtual channels hard?
+The intermediary(s) involved in the virtual channel construction additionally lock coins in virtual channels apart from the ones locked in the underlying payment channel.
+Currently, nodes get paid routing fees for every transaction they process.
+In the case of virtual channels, having a well-defined fee structure is difficult due to the following reasons: 1) The fee structure should take into account the amount of funds and the time for which these funds of the intermediary(ies) are locked in a virtual channel.
+2) It also needs to take into account the routing fee an intermediary could have earned by not locking up the coins in the virtual channel.
+RQ6: Why is off-chain dispute resolution in virtual channels hard?
+There is no offchain consensus mechanism for dispute resolution in a PCN, as opposed to the 51% honest majority assumption that exists among validators on the blockchain.
+This honest majority helps resolve disputes in the transactions posted to the blockchain.
+Designing such a dispute mechanism for layer-2 transactions is hard since transactions are private (not posted to the blockchain), and nodes do not broadcast their activities to the entire network.
+RQ7: Why is providing support for a multihop virtual channel a hard problem?
+This is hard since a multihop virtual channel construction should ensure that neither sender/receiver nor the intermediate nodes should lock coins in multiple channels at the same time.
+RQ8: Why is ensuring privacy in a virtual channel protocol a hard problem?
+In a recursive virtual channel, new virtual channels are constructed on top of existing virtual channels to facilitate payments.
+This staggered nature makes it mandatory to reveal the identity of at least one endpoint node (sender or receiver).
+This is because, at least one node among the sender/receiver is involved in all virtual channels.
+The solution to this problem is to design a multihop virtual channel.
+RQ9: Why is designing a payment channel hub (PCH) that is resistant to privacy against aborts and dynamic corruption a hard challenge?
+PCHs usually use transaction mixing for enhancing privacy, which is a process in which multiple payments from different users are mixed together in such a way that it is infeasible for the hub to link the sender and recipient of a specific transaction.
+This process helps obscure the flow of funds, providing unlinkability.
+Designing a payment channel hub that is resistant to privacy against aborts is hard because, if a PCH selectively aborts a payment from a sender/receiver, the counter party whose payment also failed can be linked.
+If sender/receiver gets corrupted during the PCH’s execution, the corresponding transaction has to be aborted to ensure atomicity, which is why the existing tumbler constructions assume a static adversary, in which certain nodes are designated as corrupted before the PCH begins execution.
+The trade off here is preserving transaction unlikability during a corrupted party’s transaction abort.
+RQ10: Why is having an offchain dispute resolution for a state channel hard?
+Current state channel protocols use an onchain transaction or a function call to the onchain smart contract in the case of dispute resolution.
 An ideal state channel protocol should be able to facilitate dispute resolutions in an offchain manner.
@@
 Onchain transactions on the hand are publicly accessible.
-VIII Network Statistics
-(a) Minimum accepted value of HTLC payment (msat)
-(b) Distribution of base fee (msat)
-(c) Distribution of rate fee (msat)
-(d) Distribution of node degree
-(e) Distribution of HTLC expiration time
-Figure 5: $\mathsf{LN}$ statistics
-This section introduces a general analysis of the parameters used in $\mathsf{LN}$, extracted from $\mathsf{LN}$ gossip messages [94].
-It offers views into user behavior and the implications of such behavior for network performance.
-Our analysis shows that the nodes prefer to choose the default value of the $\mathsf{LN}$ parameters.
-Using standard default values for these parameters allows nodes to operate smoothly together, reducing the chance of compatibility issues and enhancing the overall reliability and efficiency of the $\mathsf{LN}$.
-We evaluated the analysis on a snapshot of $\mathsf{LN}$ topology from November 27, 2021 [95, 96].
-The snapshot includes 18,331 nodes and 80,918 channels with a mean channel capacity of 4056152.35 satoshi and a median channel capacity is 1000000 satoshi (this provides us information on how funds are distributed among $\mathsf{LN}$ channels).
-Although not explicitly shown in this chart, nodes with multiple connections suggest the presence of fewer central nodes compared to the end-users nodes.
-Figure 5(a) shows that the majority of the nodes (87.9%) set the ’htlc_minimum_msat’ to a default value of 1000 msat, which indicates that the $\mathsf{LN}$ supports micropayment transactions.
-Figure 5(b) shows that ’fee_base_msat’ parameter across all of the nodes has the highest two values 1000 msat as a default value, and 0 msat as the minimum value, which indicates that most of the nodes do not set the fee_base_msat beyond the default configuration, and likely to support the network usage.
-Figure 5(c) shows the majority of the nodes (51.8%) in $\mathsf{LN}$ use various ’fee_rate_milli_msat’ values.
-A substantial portion of nodes (35.2%) set their fee rate to 1 msat to attract more transactions.
-Also, different nodes have different fee rate strategies, making $\mathsf{LN}$ a competitive environment.
-Figure 5(d) shows that the majority of the nodes (49.5%) have only one connection that indicates that in $\mathsf{LN}$, many end-users connected to fewer well-connected nodes.
-Moreover, 30% of the nodes have no payment channels, these could be new nodes added to the $\mathsf{LN}$ or they could be inactive nodes.
-Figure 5(e) shows that (81.5%) of nodes prefer to set the ’cltv_expiry_delta’ to the default value of 40 blocks, which is the number of the blocks a node can wait before losing its satoshis from $\mathsf{LN}$.
-Moreover, it helps to increase the flexibility of the $\mathsf{LN}$ to include more nodes along the payment path before reaching the ”locktime max’ parameter value (2016 blocks).
-According to the analysis, the nodes generally prefer to use the default parameter values of the lightning network.
-Using these default values can help in developing more effective routing protocols and optimizing routing fee mechanisms for the nodes.
-IX Conclusion
-In this paper, we qualitatively compared the recent work in various areas of PCN research: pathfinding and routing, virtual channels, payment channel hubs, state channels and rebalancing.
-We have also provided the gaps in research in these areas along with reasons why fulfilling those gaps is a hard research challenge.
-We hope that this paper motivates researchers to build robust protocols that address these gaps that would go a long way towards building out and developing the decentralized finance ecosystem.
+VIII Conclusion
+In this paper, we qualitatively compared the recent work in various foundational areas of PCN research: pathfinding and routing, virtual channels, state channels, payment channel hubs, and rebalancing protocols.
+We also discussed the gaps in research in these areas along with reasons why fulfilling those gaps is non-trivial.
+We hope that this paper motivates researchers to build robust protocols that address these gaps that would go a long way towards building out and developing a decentralized financial ecosystem.
+Acknowledgement
+This material is based upon work supported by the National Science Foundation under Award No.
+2148358, 1914635, 2417062, and the Department of Energy under Award No.
+DE-SC0023392.
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation and the Department of Energy.